### v1

# Elimination Distance to Dominated Clusters

Abstract
In the Dominated Cluster Deletion problem we are given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether there exists a set of at most $k$ vertices whose removal results in a graph in which each connected component has a dominating set of size at most $d$. In the Elimination Distance to Dominated Clusters problem we are again given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether we can recursively delete vertices up to depth $k$ such that each remaining connected component has a dominating set of size at most $d$. Bentert et al. [Bentert et al., MFCS 2024] recently provided an almost complete classification of the parameterized complexity of Dominated Cluster Deletion with respect to the parameters $k$, $d$, $c$, and $\Delta$, where $c$ and $\Delta$ are the degeneracy, and the maximum degree of the input graph, respectively. In particular, they provided a non-uniform algorithm with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$. They left as an open problem whether the problem is fixed-parameter tractable with respect to the parameter $k+d+c$. We provide a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters. We furthermore show that both problems are FPT when parameterized by $k+d+\ell$, where $\ell$ is the semi-ladder index of the input graph, a parameter that is upper bounded and may be much smaller than the degeneracy $c$, positively answering the open question of Bentert et al. We almost complete the picture by providing an almost full classification for the parameterized complexity and kernelization complexity of Elimination Distance to Dominated Clusters. The one difficult base case that remains open is whether Treedepth (the case $d=0$) is NP-hard on graphs of bounded maximum degree.
keywords:
Graph theory, Fixed-parameter algorithms, Dominated cluster deletion, Elimination distance
category:
[…]
1 Introduction
[…]
A related measure is the elimination distance to a class $\mathscr{C}$, which measures the number of recursive deletions of vertices needed for a graph $G$ to become a member of $\mathscr{C}$. Formally, a graph $G$ has elimination distance 00 to $\mathscr{C}$ if $G\in\mathscr{C}$, and otherwise elimination distance $k+1$, if in every connected component of $G$, we can delete a vertex such that the resulting graph has elimination distance $k$ to $\mathscr{C}$. Note that elimination distance to the class containing only the empty graph corresponds to the well-known measure treedepth. Elimination distance was introduced by Bulian and Dawar [9] in their study of the parameterized complexity of the graph isomorphism problem. Just as distance to triviality, but with more general applicability, elimination distance allows to lift algorithmic results from a base class to more general graph classes. For example, Bulian and Dawar [9] provided an FPT algorithm for the graph isomorphism problem parameterized by the elimination distance to the class $\mathscr{C}_{d}$ of graphs with maximum degree bounded by $d$, for any fixed integer $d$. Hols et al. [27] proved the existence of polynomial kernels for the vertex cover problem parameterized by the size of a deletion set to graphs of bounded elimination distance to different classes of graphs. Agrawal et al. [1] proved fixed-parameter tractability of Elimination Distance to $\mathscr{C}$ for any class $\mathscr{C}$ that can be defined by a finite set of forbidden induced subgraphs. Agrawal and Ramanujan studied the Elimination Distance to Cluster Graphs [2]. Fomin, Golovach and Thilikos further generalized several of these prior results by considering elimination distances to graph classes expressible by restricted first-order formulas [22]. These results are also implied by the recent algorithmic meta theorem for separator logic [31, 33]. The related concept of recursive backdoors has recently been studied in the context of efficient SAT solving [17, 30].
In this work, we answer the open questions of Bentert et al. [3] for Dominated Cluster Deletion and furthermore provide an almost full classification of the parameterized complexity and kernelization complexity of Elimination Distance to Dominated Clusters. More precisely, we prove that Dominated Cluster Deletion admits a uniform algorithm with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$ as well as a uniform fixed-parameter algorithm with running time $f(k,d,\ell)\cdot n^{\mathcal{O}(1)}$, where $\ell$ is the semi-ladder index of the input graph. The semi-ladder index was introduced by Fabiański et al. [19] in the study of domination-type problems and is a parameter that is upper bounded and may be much smaller than the degeneracy $c$ of a graph. In fact, our result holds for all classes of graphs where the related Partial Domination problem is fixed-parameter tractable and satisfies an additional technical condition (the concept of Partial Domination and the details will be discussed in a moment). In particular, our result answers the question of Bentert et al. positively.
Our main contribution are the uniform FPT algorithms for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on semi-ladder-free graphs.
Theorem 1.1.
Dominated Cluster Deletion can be solved in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ and in time $f(k,d,\ell)\cdot n^{\mathcal{O}(1)}$ for a computable function $f$, where $\ell$ is the semi-ladder index of the input graph.
Theorem 1.2.
Elimination Distance to Dominated Clusters can be solved in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ and $f(k,d,\ell)\cdot n^{\mathcal{O}(1)}$ for a computable function $f$, where $\ell$ is the semi-ladder index of the input graph.
The hardness results completing the classification are simple observations. We prove the following. Elimination Distance to Dominated Clusters is
  1. 1.
    para-NP-hard for parameter $k+\Delta$; this is a simple consequence of the fact that the case $k=0$ corresponds to the Dominating Set problem, which is known to be NP-hard on graphs of maximum degree $3$ [23], and
  2. 2.
    para-NP-hard for parameter $d$; this is a simple consequence of the fact that the case $d=0$ corresponds to the TreeDepth problem, which is known to be NP-hard [32], and
  3. 3.
    $\textsf{W}[2]$-hard for parameter $k+d$; this is again a simple consequence of the fact that the case $k=0$ corresponds to Dominating Set, which is W[2]-hard with respect to parameter $d$ [14], and
  4. 4.
    uniformly fixed-parameter tractable for parameter $k+d+\ell$. This is our main contribution, we discuss the details below, and
  5. 5.
    uniformly fixed-parameter tractable with respect to $k$ when $d$ is considered to be a constant, that is, admits a uniform algorithm with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$.
  6. 6.
[…]
The one case that we could not resolve is the parameter $d+\Delta$. The case $d=0$ corresponds to Treedepth, but we could not resolve whether the problem is NP-hard on graphs of bounded maximum degree. We state the following two conjectures.
Conjecture 1.1.
Treedepth is NP-hard on some class of graphs with bounded maximum degree.
Conjecture 1.2.
Treedepth is NP-hard on planar graphs.
The latter conjecture is related to the long-standing open question about the complexity of Treewidth on planar graphs. On the other hand, it is known since 1997 that Treewidth is NP-hard on graphs with maximum degree 9 [6] and more recently on graphs with maximum degree 3 [5].
1.1 Techniques
Our main contribution is the uniform algorithm with respect to the parameter $k+d+\ell$ (the algorithm for $k+d$ is a simple modification of this algorithm).
We follow the approach of Bentert et al. [3] to reduce to unbreakable graphs, that is, graphs that cannot be separated into multiple large connected components by small separators. Let us give some formal definitions. A separation in a graph $G$ is a pair of vertex subsets $A,B\subseteq V(G)$ such that $A\cup B=V(G)$ and there is no edge with one endpoint in $A\setminus B$ and the other in $B\setminus A$. The order of the separation $(A,B)$ is the size of its separator $A\cap B$. For $q,k\in\mathbb{N}$, a set $X\subseteq V(G)$ is $(q,k)$-unbreakable if for every separation $(A,B)$ of $G$ of order at most $k$, we have $|A\cap X|\leq q$ or $|B\cap X|\leq q.$ Intuitively speaking, $X$ is $(q,k)$-unbreakable if no separation of order $k$ can break $X$ in a balanced way: one of the sides must contain at most $q$ vertices of $X$. For example, cliques and $k+1$-connected graphs are $(k,k)$-unbreakable, and square grids are $(\mathcal{O}(k^{2}),k)$-unbreakable.
Lokshtanov et al. [29] proved a powerful meta theorem for logically defined graph properties. Whenever a CMSO-property (monadic second-order logic with modulo counting predicates) can be solved efficiently on unbreakable graphs (for appropriately chosen parameters $k$ and $q$), then it can also be solved efficiently on general graphs. There is a small caveat though. The meta theorem uses the non-constructive recursive understanding technique, and hence, one only obtains an efficient non-uniform algorithm on general graphs. Both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are CMSO-properties, and hence fall into this framework. Bentert et al. rely on this framework and show how to solve Dominated Cluster Deletion on unbreakable graphs in time $f(k,d)\cdot n^{\mathcal{O}(d)}$. The key combinatorial observation is that in $(q,k)$-unbreakable graphs, by the deletion of at most $k$ vertices leads to one giant component and a bounded number of remaining small components. One can then guess a dominating set for the giant component in time $\mathcal{O}(n^{d})$ and argue combinatorially to efficiently find the vertices to be deleted. Using the result of Lokshtanov et al. one obtains the non-uniform algorithm for general graphs with the desired running time.
We remark that Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are in fact a first-order definable problems, that is, for all parameters $k,d$ there exist first-order formulas of length $\mathcal{O}(k,d)$ defining the problems. Hence, the problems are fixed-parameter tractable on all classes on which first-order model checking is fixed-parameter tractable, e.g. on nowhere dense classes [24] and even monadically stable graph classes [15, 16], as well as on classes with bounded cliquewidth [10] (and classes with bounded twinwidth [8] when contraction sequences are given with the input). Note that first-order logic cannot define connected components in general, however, the connected components arising in the recursive deletion of elements must have small diameter when they can be dominated by few vertices, hence, connected components in positive instances of Dominated Cluster Deletion and Elimination Distance to Dominated Clusters become first-order definable. Thus, on these classes one obtains algorithms with running time $f(k,d)\cdot n^{\mathcal{O}(1)}$ by the meta theorems for first-order logic. It is conjectured that monadically dependent classes are the most general hereditary graph classes that admit efficient first-order model checking, we hence included them in the inclusion diagram below.
This leaves two main questions. First, can we turn the non-uniform algorithm of Bentert et al. running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ into a uniform algorithm? Second, can we improve the running time on further restricted graph classes? We answer both questions positively.
As mentioned, we follow the approach of Bentert et al. [3] to reduce to unbreakable graphs. We first observe that on unbreakable graphs, the notion of elimination distance almost collapses to distance to triviality, which allows us to treat Dominated Cluster Deletion and Elimination Distance to Dominated Clusters in a similar way. In order to avoid the non-uniform approach of Lokshtanov et al., we use a decomposition theorem due to Cygan et al. [12]. The theorem states that every graph admits a tree decomposition with small adhesion such that each bag is unbreakable in the subgraph induced by the bags below. We show that if an annotated version of the Partial Dominating Set problem can be solved in a slightly modified graph (that we call the bag graph), then we can implement an efficient dynamic programming algorithm along the tree decomposition to the original input graph for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters.
The (Annotated) Partial Dominating Set problem is the following problem. We are given a graph with three sets $F$, $R$ and $B$ and parameters $d$ and $k$. We ask whether there exist $k$ vertices of $G-F$ ($F$ stands for forbidden) that may be deleted such that the remaining graph has a Red-Blue dominating set of size at most $d$, i.e., the vertices of $R$ can be dominated by $d$ vertices of $B$ ($R$ and $B$ stand for Red and Blue). Formally, the problem is defined as follows. Given a graph $G$ with unary predicates $F\subseteq V(G)$, $R\subseteq V(G)$, and $B\subseteq V(G)$ and parameters $d$ and $k$, does there exist a set $D\subseteq B$ of size at most $d$ and a set $X\subseteq V(G)-F$ of size at most $k$ such that $G[R-X]$ is dominated by $D$? The problem can obviously be solved in time $n^{\mathcal{O}(d)}$ (guess the dominating set and check if at most $k$ Red vertices remain undominated). By our dynamic programming approach we hence get a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ in general graphs. We then turn our attention to restricted graph classes where (Annotated) Partial Domination can be solved more efficiently. Using an approach of Fabiański et al. [19] we observe that (Annotated) Partial Domination can be solved efficiently on all semi-ladder-free graph classes, which yields our second main result.
Figure 1: Relation between the discussed graph classes. Arrows indicate inclusion. Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are fixed-parameter tractable on monadically stable classes by a meta theorem for first-order logic. The same is true for bounded twinwidth classes when contraction sequences are given with the input. Our main result is the fixed-parameter tractability on semi-ladder-free graphs.
Semi-ladder-free are very general graph classes. For example, all degenerate classes are biclique-free (that is, exclude a fixed complete bipartite graph $K_{t,t}$ as a subgraph), which in turn are semi-ladder-free, hence, our result in particular answers the question of Bentert et al. positively. On the other hand, these classes are incomparable with monadically stable classes and classes with bounded twinwidth on which we can efficiently solve the first-order model checking problem. For example, the class of all ladders has bounded twinwidth (in fact even bounded linear cliquewidth) but is not semi-ladder-free. The class of all co-matchings is monadically stable (and has bounded linear cliquewidth as well) and is not semi-ladder-free. On the other hand, the class of all $1$-subdivided cliques is semi-ladder-free (and even $2$-degenerate) but is neither monadically stable nor has bounded twinwidth.
As mentioned, we require that (Annotated) Partial Domination needs to be solved on a slightly modified graph class (on the bag graphs), however, the property of being semi-ladder-free is preserved by this modification, so that ultimately we conclude our main results, Theorem 1.1 and Theorem 1.2.
Organization. The paper is organized as follows. After recalling the necessary notation in Section 2, we talk more precisely about partial domination and the graph parameters enabling FPT evaluation in Section 3. We then present how to solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graph with bounded semi-ladder index in Section 4. After that, in Section 5, we sketch the uniform algorithm on general semi-ladder-free graphs for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$. The details are in appendix. The hardness results conclude the paper in Section 6.
2 Preliminaries
Graphs. We consider finite, undirected graphs without loops. Our notation is standard, and we refer to Diestel’s textbook for more background on graphs [13]. We write $|G|$ for the size of the vertex set $V(G)$ and $\|G\|$ for $|V(G)|+|E(G)|$, where $E(G)$ is the edge set of a graph $G$. A connected component of $G$ is a maximal connected subgraph of $G$. For a vertex subset $X\subseteq V(G)$ we write $G[X]$ for the subgraph of $G$ induced by $X$. We write $G-X$ for $G[V(G)\setminus X]$ and for singleton sets $\{v\}$ we write $G-v$ instead of $G-\{v\}$. We denote the open neighborhood of a vertex $v$ by $N(v)$ and the closed neighborhood (including the vertex $v$) by $N[v]$. For a set $Y\subseteq V(G)$ we write $N[Y]$ for $\bigcup_{v\in Y}N[y]$.
Semi-ladders. Let $G$ be a graph. Two sequences, $a_{1},\ldots,a_{n}\in V(G)$ and $b_{1},\ldots,b_{n}\in V(G)$ of $2n$ distinct vertices, form a semi-ladder of order $n$ in $G$ if $\{a_{i},b_{j}\}\in E(G)$ for all $i,j\in\{1,\ldots,n\}$ with $i>j$, and $\{a_{i},b_{i}\}\notin E(G)$ for all $i\in\{1,\ldots,n\}$. Note that we do not impose any condition for $i<j$. The semi-ladder index of a graph is the maximum order of a semi-ladder that it contains. A class of graphs has bounded semi-ladder index or is called semi-ladder-free if there is some $\ell$ such that the semi-ladder index of all of its members is bounded by $\ell$, see Figure 2.
$a_{1}$$a_{2}$$a_{3}$$a_{4}$$b_{1}$$b_{2}$$b_{3}$$b_{4}$
Figure 2: Semi-ladder of order 4. Solid lines represent edges while dashed edges represent non-edges.
Tree decompositions. A tree decomposition of a graph $G$ is a pair $\mathcal{T}=(T,\mathsf{bag})$, where $T$ is a rooted tree and $\mathsf{bag}\colon V(T)\to 2^{V(G)}$ is a mapping that assigns to each node $x\in T$ its bag $\mathsf{bag}(x)\subseteq V(G)$ such that the following conditions are satisfied:
  • For every vertex $v\in V(G)$, there exists a node $x\in V(T)$ such that $v\in\mathsf{bag}(x)$.
  • For every vertex $v\in V(G)$, the set of nodes $x\in V(T)$ satisfying $v\in\mathsf{bag}(x)$ induces a connected subtree of $T$.
  • For every edge $\{u,v\}\in E(G)$, there exists a node $x\in V(T)$ such that $u,v\in\mathsf{bag}(x)$.
Let $\mathcal{T}=(T,\mathsf{bag})$ be a tree decomposition of a graph $G$. The adhesion of a node $x\in V(T)$ is defined as $\mathsf{adh}(x)\coloneqq\mathsf{bag}(\mathsf{parent}(x))\cap\mathsf{bag}(x)$ and the margin of a node $x\in V(T)$ is defined as $\mathsf{mrg}(x)\coloneqq\mathsf{bag}(x)\setminus\mathsf{adh}(x)$. The cone at a node $x\in V(T)$ is defined as $\mathsf{cone}(x)\coloneqq\bigcup_{y\succcurlyeq_{T}x}\mathsf{bag}(y)$ and the component at a node $x\in V(T)$ is defined as $\mathsf{comp}(x)\coloneqq\mathsf{cone}(x)\setminus\mathsf{adh}(x)=\bigcup_{y\succcurlyeq_{T}x}\mathsf{mrg}(y)$. Here, $y\succcurlyeq_{T}x$ means that $y$ is a descendant of $x$ in $T$.
Unbreakable graphs. A separation in a graph $G$ is a pair of vertex subsets $A,B\subseteq V(G)$ such that $A\cup B=V(G)$ and there is no edge with one endpoint in $A\setminus B$ and the other in $B\setminus A$. The order of the separation $(A,B)$ is the size of its separator $A\cap B$.
For $q,k\in\mathbb{N}$, a vertex subset $X$ in a graph $G$ is $(q,k)$-unbreakable if for every separation $(A,B)$ of $G$ of order at most $k$, we have $|A\cap X|\leq q$ or $|B\cap X|\leq q.$ Not every graph is $(q,k)$-unbreakable, but all graphs admit tree decompositions with small adhesion and unbreakable parts. For fixed $q,k\in\mathbb{N}$, a tree decomposition $(T,\mathsf{bag})$ of a graph $G$ is $(q,k)$-unbreakable if for every node $x\in V(T)$, the bag $\mathsf{bag}(x)$ is $(q,k)$-unbreakable in $G[\mathsf{cone}(x)]$. By the result of Cygan et al. [12], such tree decompositions exist and can be computed efficiently.
Theorem 2.1 (Theorem 10 of [12]).
Let $G$ be a graph and $k$ an integer. There exists an integer $q\in 2^{\mathcal{O}(k^{2})}$ and an algorithm that computes in time $q\cdot|G|^{2}\|G\|$ a tree decomposition $(T,\mathsf{bag})$ of $G$ with at most $|G|$ nodes such that the following conditions hold:
  • For every node $x\in V(T)$, the bag $\mathsf{bag}(x)$ is $(q,k)$-unbreakable in the subgraph induced by $G[\mathsf{cone}(x)]$.
  • For every node $x\in V(T)$, the adhesion $\mathsf{adh}(x)$ is of size at most $q$.
A tree decomposition $(T,\mathsf{bag})$ is regular if for every non-root node $x\in V(T)$,
  • the margin $\mathsf{mrg}(x)$ is non-empty,
  • the graph $G[\mathsf{comp}(x)]$ is connected, and
  • every vertex of $\mathsf{adh}(x)$ has a neighbor in $\mathsf{comp}(x)$.
We may further assume that the tree decomposition we get from Theorem 2.1 is regular, see the construction of Lemma 2.8 of [7] and the discussion in [31].
Parameterized complexity. A parameterized problem $L\subseteq\Sigma^{*}\times\mathbb{N}$ is fixed-parameter tractable (FPT) if there exists an algorithm that decides for an instance $(x,k)$ whether $(x,k)\in L$ in time $f(k)\cdot|(x,k)|^{\mathcal{O}(1)}$ for some computable function $f$.
3 Partial domination and FPT evaluation
We show how to efficiently solve Annotated Partial Domination problem on semi-ladder-free graphs. This problem falls into the framework of domination type problems on semi-ladder-free graphs introduced by Fabiański et al. [19].
We are going to use the following meta theorem of Fabiański et al. [19]. A domination-type problem is a formula $\exists\bar{x}\forall\bar{y}\delta(\bar{x},\bar{y})$, where $\delta$ is a positive first-order formula using fixed distances as predicates (called a distance formula in [19]). On every graph $G$ a formula $\delta(\bar{x},\bar{y})$ defines a bipartite graph $G_{\delta}$ on the vertex $V(G)^{|\bar{x}|}\times V(G)^{|\bar{y}|}$, where two elements $\bar{u}$ and $\bar{v}$ are connected by an edge if the formula $\delta(\bar{u},\bar{v})$ is true in $G$. The semi-ladder index of $\delta$ on $G$ is defined as the semi-ladder index of $G_{\delta}$.
Theorem 3.1 (Theorem 19 of [19]).
Let $\mathscr{C}$ be a class of graphs and let $\delta$ be a distance formula. Assume that $G_{\delta}$ has bounded semi-ladder index $\ell$ for some fixed $\ell\in\mathbb{N}$ on all graphs $G\in\mathscr{C}$. Then there is an algorithm that solves the domination-type problem $\delta$ on graphs $G$ from $\mathscr{C}$ in time $f(\ell,|\varphi|)\cdot\|G\|$.
Annotated Partial Domination can be written with the following first order formula:
$\varphi(k,d)=\exists x_{1}\ldots\exists x_{k}$ $\exists y_{1}\ldots\exists y_{d}\forall z\leavevmode\nobreak\ \leavevmode\nobreak\ \Big{(}\bigwedge_{1\leq i\leq k}\neg F(x_{i})\Big{)}\wedge\Big{(}\bigwedge_{1\leq i\leq d}B(y_{i})\Big{)}\wedge$
$\Big{(}R(z)\rightarrow\Big{(}\bigvee_{1\leq i\leq k}\mathrm{dist}(z,x_{i})=0\vee\bigvee_{1\leq i\leq d}\mathrm{dist}(z,y_{i})\leq 1\Big{)}\Big{)}$
However, $\varphi$ is not a distance formula as it uses negations. This can be circumvented by inverting the predicate $F^{\prime}=V(G)-F$ and $R^{\prime}=V(G)-R$. We then obtain the following formula
$\delta_{k,d}=\exists x_{1}\ldots\exists x_{k}$ $\exists y_{1}\ldots\exists y_{d}\forall z\leavevmode\nobreak\ \leavevmode\nobreak\ \Big{(}\bigwedge_{1\leq i\leq k}F^{\prime}(x_{i})\Big{)}\wedge\Big{(}\bigwedge_{1\leq i\leq d}B(y_{i})\Big{)}\wedge$
$\Big{(}R^{\prime}(z)\vee\bigvee_{1\leq i\leq k}\mathrm{dist}(z,x_{i})=0\vee\bigvee_{1\leq i\leq d}\mathrm{dist}(z,y_{i})\leq 1\Big{)}$
We prove next that $\delta_{d,k}$ has bounded semi-ladder index on every graph $G$ with bounded semi-ladder index.
Lemma 3.2.
Let $G$ be a graph with semi-ladder index $\ell$. Then $\delta_{d,k}$ has semi-ladder index at most $f(d,k,\ell)$ on $G$, for some computable function $f$.
Proof 3.3.
Lemma 4 of [19] states that if $\varphi_{1}(\bar{x};\bar{y}),\ldots,\varphi_{k}(\bar{x};\bar{y})$ are formulas and $\psi(\bar{x};\bar{y})$ is a positive boolean combination of $\varphi_{1},\ldots,\varphi_{k}$, and $G$ is a graph such that $\varphi_{1}(G),\ldots,\varphi_{k}(G)$ have bounded semi-ladder index, then also $\psi(G)$ has bounded semi-ladder index. This is the case for the above formula $\delta_{k,d}$. By Lemma 5 of [19], the formula $\bigvee_{1\leq i\leq d}\mathrm{dist}(z,y_{i})\leq 1$ has bounded semi-ladder index on every graph with bounded semi-ladder index. All other formulas trivially have bounded semi-ladder index on all graphs.
Corollary 3.4.
Let $\mathscr{C}$ be a class of graphs with bounded semi-ladder index $\ell$. Then Annotated Partial Domination can be solved in time $f(d,k,\ell)\cdot\|G\|$ for all $G\in\mathscr{C}$ for a computable function $f$.
While the semi-ladder index is currently one of the most general graph parameter enabling FPT evaluation of $\delta_{k,d}$, some other parameter could have this property.
Property 1.
A graph parameter $t(\cdot)$ has 1 if for every graph $G$ Annotated Partial Domination can be solved in time $f(d,k,t(G))\cdot|G|^{\mathcal{O}(1)}$ for a computable function $f$.
As previously discussed, examples of parameters with 1 include treewidth, maximum degree, Hadwiger number (i.e. largest clique minor), degeneracy, largest excluded bi-clique, as they all imply bounded semi-ladder index. Other parameters, which are however not as interesting, as the problem can be solved by efficient FO model checking, are cliquewidth and twinwidth (when contraction sequences are given with the input).
As mentioned, we want to solve Annotated Partial Domination on slightly modified graphs. See Definition B.6 for the precise definition of bag graphs. We therefore want a parameter that, if bounded on the graph, bounds a parameter with 1 on bag graphs.
Property 2.
A graph parameter $t_{2}(\cdot)$ has 2 if there is a parameter $t_{1}(\cdot)$ satisfying 1 such that for every graph $G$, and integers $q,k,d$, every $(q,k)$-unbreakable tree decomposition $(T,\mathsf{bag})$, vertex $x\in T$, and set $S\subseteq\mathsf{cone}(x)$ we have that $t_{1}(\mathsf{bgraph}_{S}(x))\leq f(q,k,d,t_{2}(G))$.
In the rest of the paper, we precisely reference parameters having 1 or 2. In particular, the semi-ladder index also satisfies 2, see Lemma B.11.
4 Algorithm on unbreakable graphs based on Annotated Partial Domination
As sketched above, we first solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graphs. Our algorithm runs efficiently on all graph classes for which we can efficiently solve the Annotated Partial Domination problem, that is, for classes on which a parameter $t(\cdot)$ satisfying 1 is bounded.
In the following, $q,k$, and $d$ will always be non-negative integers, and we will not always quantify them for readability. We will also tacitly assume that our graphs have at least $12q+$ 1 vertices. Then, whenever we delete a set $S$ of at most $k$ vertices there will be a unique connected component $C_{0}$ of $G-S$ with more than $q$ vertices that we call the large connected component. When a graph has less vertices, we can solve all our problems by brute-force.
Lemma 4.1.
Let $G$ be a $(q,k)$-unbreakable graph, and let $v\in V(G)$. Then $G-v$ is $(q,k-1)$-unbreakable.
Proof 4.2.
Assume that there is a connected component of $G-v$ that is not $(q,k-1)$-unbreakable. Then, there is a separation $(A,B)$ of $G-v$ such that $|A|\geq q$, $|B|\geq q$ and $|A\cap B|\leq k-1$. By adding the vertex $v$ to the separator, we obtain a separation $(A^{\prime},B^{\prime})$ of $G$ such that $|A^{\prime}|>q$, $|B^{\prime}|>q$ and $|A^{\prime}\cap B^{\prime}|\leq k$ contracting that $G$ is $(q,k)$-unbreakable.
Lemma 4.3.
Let $G$ be a $(q,k)$-unbreakable graph. If $G$ is a positive instance of Dominated Cluster Deletion with parameters $k$ and $d$, then $G$ has a dominating set of size at most $q+d$.
Proof 4.4.
Let $S^{\prime}$ with $|S^{\prime}|\leq k$ be a solution for Dominated Cluster Deletion with parameters $d$ and $k$. Let $A=S^{\prime}\cup C_{0}$, where $C_{0}$ is the large connected component of $G-S^{\prime}$ and $B=S^{\prime}\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of $G-S^{\prime}$. Then $(A,B)$ is a separation of order at most $k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable. As $G$ is a positive instance of Dominated Cluster Deletion, $C_{0}$ can be dominated by a set $D_{0}$ of size at most $d$. Then $D_{0}\cup B$ is dominating set of $G$ of size at most $q+d$.
4.1 Dominated Cluster Deletion
We first deal with Dominated Cluster Deletion. Our approach is based on finding skeletons for dominated cluster deletion, which are vertex sets that can be extended to the deletion set of a solution for Dominated Cluster Deletion and in $(q,k)$-unbreakable graphs capture the part of a solution that essentially separates the large connected component from the other connected components.
Definition 4.5.
Let $G$ be a $(q,k)$-unbreakable graph. We call a set $S$ a skeleton for dominated cluster deletion for parameters $k$ and $d$ if there is a superset $S^{\prime}\supseteq S$ of vertices satisfying:
  • $|S^{\prime}|\leq k$,
  • every connected component of $G-S^{\prime}$ has domination number at most $d$, and
  • $S$ contains exactly the vertices of $S^{\prime}$ that have at least one neighbor in $C^{\prime}_{0}$ and at least one neighbor in $G-N[C^{\prime}_{0}]$, where $C^{\prime}_{0}$ is the large connected component of $G-S^{\prime}$.
Note that $S$ may be the empty set in the third item above if $G-S^{\prime}$ has only one connected component. Figure 3 shows examples of skeletons for dominated cluster deletion.
Figure 3: Different skeletons for the dominated cluster deletion problem for parameter $k=4$ and $d=1$ on a fixed (14,4)-unbreakable graph. The skeleton $S$ is in red, the witness $S^{\prime}$ also contains the crossed-out vertices.
The next lemma is the key ingredient to compute the set of skeletons (and to show that there are only few possible skeletons). It shows that there is a small set of vertices that hits every skeleton.
Lemma 4.6.
Let $q,k,d$ be integers, $G$ a $(q,k)$-unbreakable graph, and assume that $S$ is a non-empty skeleton for dominated cluster deletion for parameters $k$ and $d$. Given a dominating set $X$ of size at most $q+d$, there is a vertex $v\in S$ that is
  1. 1.
    in the dominating set $X$, or
  2. 2.
    in the neighborhood $N(x)$ of a vertex $x\in X$, where $\deg(x)\leq q$, or
  3. 3.
    in the neighborhood $N(y)$ of a vertex $y\in N(x)$, where $\deg(x)\leq q$, $\deg(y)\leq q$, and $x\in X$.
Proof 4.7.
Assume that the Case 1 is not satisfied. Meaning that no vertices from the skeleton are in the dominating set $X$. Denote by $C^{\prime}_{0}$ the large connected component of $G-S^{\prime}$ for a set $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, hence, $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$. Let $C_{0}$ be the large connected component of $G-S$. Note that it contains $C^{\prime}_{0}$.
Let $A=S\cup C_{0}$ and $B=S\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of $G-S$. Then $(A,B)$ is a separation of order at most $|S|\leq k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable. Pick an arbitrary element $v\in S$. By definition of $S$, $v$ has a neighbor in $C_{0}$ and in a second connected component, say $C_{1}$ of $G-S$. This neighbor must be a vertex $w$ of degree at most $q$ as its neighborhood lies in $G-C_{0}\subseteq B$. If $w$ is in $X$, then this vertex $v$ satisfies Case 2. Otherwise, $w$ must have a neighbor $x$ in $X$ (since $X$ is a dominating set). Since we are not in Case 1, this $x$ is not in $S$, and is therefore in $C_{1}$. Hence, similarly to $w$, we have that $x$ has degree at most $q$, and we are in Case 3.
Lemma 4.6 gives rise to a bounded search tree branching algorithm that branches over the possible choices for a vertex from $S$ given by the conditions of the lemma.
Definition 4.8.
The skeleton-algorithm with parameter $q,k,d$ over a graph $G$ returns a family of sets of vertices, each set of size at most $k$. First, it computes a dominating set $X$ of size at most $q+d$ (returning the empty family if there is no such set). Second, it computes the set $X\cup Y\cup Z$, where $Y$ is the set of all vertices in the neighborhood $N(x)$ of a vertex $x\in X$ with $\deg(x)\leq q$, and the set $Z$ is the set of all vertices in the neighborhood $N(y)$ of a vertex $y\in Y$ where $\deg(y)\leq q$. Third, call the algorithm with parameters $q,k-1,d$ on $G-v$ for every $v$ in $X\cup Y\cup Z$ (and get a family $\mathcal{F}_{v}$). Finally, for every $v$ in $X\cup Y\cup Z$ add $v$ to every set in $\mathcal{F}_{v}$, and return the empty set and the union of all the families $\mathcal{F}_{v}$.
Note that in the above definition, if a family $\mathcal{F}_{v}$ is empty, adding $v$ to every set still results in an empty family (e.g. in the case where $G-v$ does not have a $q+d$ dominating set). To not be confused with $\mathcal{F}_{v}$ having the empty set as one of the family member (e.g. for branches of the algorithm that stops before reaching $k=0$).
Algorithm 1 Skeleton-algorithm
1:procedure Skeleton($G,q,k,d$)
2:     $\mathcal{F}\coloneqq\emptyset$
3:     if $k=0$ then
4:         return $\mathcal{F}$
5:     end if
6:     $X\coloneqq$ DominatingSet($G,q+d$)
7:     $Y\coloneqq\{N(x)\colon x\in X\wedge\deg(x)\leq q\}$
8:     $Z\coloneqq\{N(y)\colon y\in Y\wedge\deg(y)\leq q\}$
9:     for all $v\in X\cup Y\cup Z$ do
[…]
11:     end for
12:     return $\mathcal{F}$
13:end procedure
We now show that the skeleton algorithm 1) runs efficiently and outputs a small family of sets. This is proved in Lemma 4.9, and 2) if the given graph was a positive instance to Dominated Cluster Deletion, then the skeleton algorithm outputs all possible skeletons; see Lemma 4.11.
Lemma 4.9.
Let $t(\cdot)$ be a graph parameter satisfying 1, then the skeleton-algorithm runs in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, and outputs a family of at most $g(q,k,d)$ sets where $f$ and $g$ are computable functions.
Proof 4.10.
Computing $X,Y$ and $Z$ takes time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$ thanks to the fact that $t(\cdot)$ satisfies 1. We then branch over $(q+d)q^{2}$ many choices for $v$, and recurse at most $k$ times.
Lemma 4.11.
If $G$ is a $(q,k)$-unbreakable graph that is a positive instance to Dominated Cluster Deletion with parameters $(k,d)$, then the skeleton-algorithm outputs a family that contains every skeleton.
Proof 4.12.
By induction on $k$. If $k=0$, we simply output the empty set. Otherwise, by Lemma 4.3 there is such a set $X$, and by Lemma 4.6 at least one element $v$ of $X$ is in $S$. For the correct choices of $v$, $G-v$ is a positive instance for Dominated Cluster Deletion with parameter $k-1,d$, and by Lemma 4.1, $G-v$ is $(q,k-1)$-unbreakable, so the recursive call to Skeleton produces the expected sets by induction.
It remains to verify that $S$ can indeed be enlarged to a set $S^{\prime}$ of size at most $k$ with the properties.
Lemma 4.13.
Given a $(q,k)$-unbreakable graph $G$ and a set $S$ of at most $k$ vertices, we can test in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, whether $S$ is a skeleton for Dominated Cluster Deletion with parameters $k$ and $d$, and $t(\cdot)$ satisfies 1.
Proof 4.14.
We need to check the existence of a set $S^{\prime}\supseteq S$ such that:
  • every connected component of $G-S^{\prime}$ has domination number at most $d$, and
  • $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$ where $C^{\prime}_{0}$ is the large connected component of $G-S^{\prime}$.
Note that by these conditions the large connected component $C_{0}$ of $G-S$ is separated from the small connected components by $S$. What remains to be done is hence the following.
  1. 1.
    Find an optimal solution for $G-S-C_{0}$. This can be done in time $f(k,d)$, because this subgraph has size at most $q$ and yields a value $k^{\prime\prime}$ of vertices that need to be deleted. If $k^{\prime\prime}>k-k^{\prime}$ we reject the instance (in this branch).
  2. 2.
    Test whether an additional set $W$ of $k^{\prime\prime\prime}=k-k^{\prime}-k^{\prime\prime}$ vertices can be deleted from $C_{0}$ such that $G[C_{0}-W]$ can be dominated by $d$ vertices. Note that by assumption on $S$ and $S^{\prime}$ the deletion of $W$ will not cause $C_{0}$ to break into multiple further connected components. Hence, the remaining problem is an instance of Partial Domination with parameter $k^{\prime\prime\prime}$. If $S$ is a skeleton, such a set exists (as $S^{\prime}\cap C_{0}$ is a candidate) and hence, if $G$ is a positive instance of Dominated Cluster Deletion a positive solution will be found. Since $t(\cdot)$ is a graph parameter satisfying 1, we can test the existence of such $W$ in time $f(k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$. If no solution is found we reject the instance (in this branch).
The combination of Lemmas 4.6, 4.9, 4.11 and 4.13 yields the following theorem.
Theorem 4.15.
Dominated Cluster Deletion on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
4.2 Elimination Distance to Dominated Clusters
We now show how to adapt the algorithm for Elimination Distance to Dominated Clusters. The approach is very similar. Of course, the notion of skeletons has to be adapted, however, they behave very similarly on unbreakable graphs.
First let’s properly define Elimination Distance to Dominated Clusters.
Definition 4.16.
A graph $G$ has elimination distance $k$ to a class $\mathscr{C}$ if:
  • $G$ is not connected and every connected component of $G$ has elimination distance $k$ to $\mathscr{C}$, or
  • $G$ is connected and there is a vertex $v$ such that $G-v$ has elimination distance $k-1$ to $\mathscr{C}$, or
  • $G\in\mathscr{C}$ and $k=0$.
In our case we measure, for every $d$, the elimination distance to the class $\mathscr{C}_{d}$ of $d$-dominated clusters i.e., graphs where every connected component has a $d$-dominating set. We interchangeably write “elimination distance $k$ to $d$-dominated cluster”, and “a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters”. For algorithmic purposes, the above definition is not the most useful. We will instead look at the set of vertices that needs to be deleted to get $d$-dominated clusters. While this set is of unbounded size, it has a nice tree structure that is much more useful for our algorithms. This characterization is pretty standard, see the definition of elimination order [9, Proposition 4.3]
For a rooted tree $T$, we write $\preccurlyeq_{T}$ for the ancestor relation, that is, $x\preccurlyeq_{T}y$ if $x$ lies on the unique path from $y$ to the root of $T$. Note that we treat every node as an ancestor of itself. The least common ancestor of two nodes $y,z\in V(T)$ is the $\preccurlyeq_{T}$-maximal element $x$ with $x\preccurlyeq_{T}y$ and $x\preccurlyeq_{T}z$.
Definition 4.17.
Let $G$ be a graph and $S\subseteq V(G)$. We say that $S$ is tree-structured if there is a rooted tree $T$ and a mapping $\lambda:S\rightarrow V(T)$ such that every path in $G$ between two vertices $u,v$ of $S$ must contain a common ancestor in $T$ of $\lambda(u)$ and $\lambda(v)$. The tree $T$ is called an elimination tree for $S$. The elimination depth of a tree-structured set with respect to the tree $T$ is the depth of $T$. The elimination depth of $S$ is the minimum elimination depth of $S$ over all elimination trees for $S$.
Remark 4.18.
A graph $G$ is a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters if and only if there is a tree-structured set $S$ with elimination depth $k$ such that every connected component of $G-S$ has a $d$-dominating set.
Lemma 4.19.
Let $G$ be a $(q,k)$-unbreakable graph. Let $S$ be a tree-structured set with an elimination tree $T$ of depth at most $k$. Then
  1. 1.
    the largest connected component $C_{0}$ of $G-S$ is uniquely determined, and
  2. 2.
    for every component $C$ of $G-S$ there is a subset $S(C)\subseteq S$ of size at most $k$ that separates $C$ from the rest, that is, $G-S(C)$ has $C$ as one of its components.
Proof 4.20.
Let $C$ be a connected component of $G-S$. As $S$ is tree-structured by $T$, $C$ is not adjacent to two $\preccurlyeq_{T}$-incomparable elements, that is, the set $S(C)$ of neighbors of $C$ in $S$ consists of $\preccurlyeq_{T}$-comparable elements. $C$ is a connected component both of $G-S$ and of $G-S(C)$, as all neighbors of $C$ in $S$ are collected in $S(C)$.
Assume now that there are two connected components $C_{0},C_{1}$ of $G-S$ of size greater than $q$. However, $C_{0}$ and $C_{1}$ are separated by $S(C_{0})$ of size at most $k$, which contradicts the $(q,k)$-unbreakability of $G$ (assuming that $G$ has at least $12q+$ 1 vertices). Hence, the largest connected component $C_{0}$ of $G-S$ is the unique connected component of size greater than $q$.
Lemma 4.21 (Analog of Lemma 4.3).
Let $G$ be a $(q,k)$-unbreakable graph. If $G$ is a positive instance of Elimination Distance to Dominated Clusters with parameters $k$ and $d$, then $G$ has a dominating set of size at most $q+d$.
The proof can be taken verbatim from Lemma 4.3 with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
Definition 4.22.
Let $G$ be a $(q,k)$-unbreakable graph. We call a set $S$ a skeleton for elimination distance to dominated clusters for parameters $k$ and $d$ if there is a tree-structured superset $S^{\prime}\supseteq S$ of elimination depth at most $k$ satisfying:
  • every connected component of $G-S^{\prime}$ has domination number at most $d$, and
  • $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$ where $C^{\prime}_{0}$ is the large connected component of $G-S^{\prime}$ (the next lemma shows that this connected component is uniquely determined even though $S^{\prime}$ can be larger than $k$).
The following observation follows immediately from Lemma 4.19. {observation} Every skeleton has size at most $k$.
Figure 4: Example of a skeleton for elimination distance to dominated clusters for parameter $k=3$ and $d=1$ on a (12,3)-unbreakable graph (left). The skeleton $S$ is in red, the witness $S^{\prime}$ also contains the crossed-out vertices. A tree structure for $S^{\prime}$ of elimination depth $3$ is on the right.
Now the analog of Lemma 4.6 also holds for skeletons for elimination distance to dominated clusters. The proof can be taken verbatim with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
Lemma 4.23.
Let $q,k,d$ be integers, $G$ be a $(q,k)$-unbreakable graph and assume that $S$ is a non-empty skeleton for Elimination Distance to Dominated Clusters with parameters $k$ and $d$. Given a domination set $X$ of size $q+d$, there is a vertex $v\in S$ that is
  1. 1.
    in the dominating set $X$, or
  2. 2.
    in the neighborhood $N(x)$ of a vertex $x\in X$ where $\deg(x)\leq q$, or
  3. 3.
    in the neighborhood $N(y)$ of a vertex $y\in N(x)$ with $\deg(x)\leq q$, $\deg(y)\leq q$, and $x\in X$.
Hence, the algorithm to guess a skeleton $S$ can be carried out in exactly the same way as for Dominated Cluster Deletion. We then guess a linear order on $S$ (which takes time at most $\mathcal{O}(k!)$). In the final step when we aim to extend a fixed skeleton $S$ to a tree-structured solution $S^{\prime}$, the only difference is that we have to solve the base problem Elimination Distance to Dominated Clusters for the small part $G-S-C^{\prime}_{0}$, which again can be simply solved by brute-force. Note that because $G-S-C^{\prime}_{0}$ is small we can also find an elimination tree $T$ for $S^{\prime}$ and embed $S$ according to the originally guessed order. We conclude the main theorem of this section. The full algorithm is presented in the appendix.
Theorem 4.24.
Elimination Distance to Dominated Clusters on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
4.3 Skeletons for general graphs
In this section, we show that by parameterizing only by $q,k$, and $d$ we can compute the skeleton of unbreakable graphs in time $f(q,k,d)\cdot|G|^{\mathcal{O}(d)}$. This construction follows the work of Bentert et al. [3]. Having access to this algorithm will then be useful for our uniform version of their result.
Lemma 4.25.
There is an algorithm that, given a $(q,k)$-unbreakable graph $G$, computes all skeletons $S$ for dominated cluster deletion and for elimination distance to dominated clusters for parameters $k$ and $d$ in time $f(q,k,d)\cdot n^{\mathcal{O}(d)}$.
This lemma resembles and can be proved analogously to [3, Lemma 7].
Proof 4.26.
Assume the existence of a solution $S^{\prime}$ and let $S$ be it skeleton. We guess a set $D$ of at most $d$ vertices in time $O(n^{d})$. (This will play the role of the dominating set for the big component). We then define $S_{1}=G-N[D]$ and $S_{2}=\{v\in N(u)\leavevmode\nobreak\ |\leavevmode\nobreak\ u\in S_{1}\wedge\deg(u)\leq q\}$. If either $|S_{1}|$ or $|S_{2}|$ is bigger than $q$, we terminate this branch. Otherwise, we now claim that $S\subseteq S_{2}$.
To this end, let $C^{\prime}_{0}$ be the large connected component of $G-S^{\prime}$ and assume that the algorithm guessed $D$ as a dominating set for $C^{\prime}_{0}$. Observe that $C^{\prime}_{0}\subseteq N[D]\subseteq N[C^{\prime}_{0}]$. Also, a vertex of $G-N[C^{\prime}_{0}]$ has no neighbor in $C^{\prime}_{0}$, and therefore only has degree at mots $q$. Finally, remember that any vertex $v$ of $S$ must have neighbor $u$ that is not in $N[C^{\prime}_{0}]$ and hence not in $N[G]$. So $u$ is in $S_{1}$ and $u$ has degree at most $q$. So $v$ belongs to $S_{2}$.
We conclude that $S\subseteq S_{2}$. So we can guess $S$ in time $f(q,k,d)\cdot n^{\mathcal{O}(d)}$.
Note that Lemma 4.25 applies to skeletons for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters. Furthermore, the dominating set $D$ for the large connected component is computed (by brute-force).
5 Annotated versions for Dominated Cluster Deletion
Definition 5.1 (Annotated Dominated Cluster Deletion).
Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Dominated Cluster Deletion problem is to find at most $k$ vertices $S\subseteq G-F$ such that every connected component $C$ of $G-S$ has a Red-Blue dominating set of size at most $d$ (i.e. $d$ Blue vertices dominating all Red vertices).
The next proposition states that this problem can be solved on specific instances which we call bag graphs.
Definition 5.2.
A graph $G$ is called a $(q,k,d)$-bag graph, if the vertices of $G$ can be partition into two vertex sets $I,E$ called respectively interior vertices and exterior vertices, such that:
  • a connected component of $G[E]$ has at most $12d+$ 1 vertices,
  • a connected component of $G[E]$ has at most $q$ neighbors in $I$, and
  • for every separation $L,R$ of $G$ of order $k$ such that $L\cap R\subseteq I$, we have that either $|L\cap I|\leq q$, or $|R\cap I|\leq q$.
Proposition 5.3.
There is an algorithm such that, given a $(q,k,d)$-bag graph $G=(I,E)$, and vertex sets $F,R,B$ of $G$ satisfying $E\subseteq F$ and $|E\cap B|\leq qd$, the algorithm decides whether $G$ is a positive instance to $(k,d)$-Annotated Dominated Cluster Deletion, in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
We prove Proposition 5.3 in Appendix A. Using this, we then prove the following statement via dynamic programming on an unbreakable tree decomposition in Appendix B.
Theorem 5.4.
There is an algorithm solving Annotated Dominated Cluster Deletion in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying 2.
Remember that 2 implies that, on the bag graphs that are build during the dynamic programming, some parameter with 1 is bounded, hence we can use Proposition 5.3 these bag graphs. Note that Theorem 5.4 implies Theorem 1.1, by setting $F=\emptyset$ and $R=B=V(G)$. The annotated versions for Elimination Distance to Dominated Clusters is introduced and solved in Appendix C.
6 Hardness Results
In this section we give the details of the hardness results mentioned in the introduction.
Para-NP-hardness for parameter $k+\Delta$ and for parameter $d$
A problem is para-NP-hard with respect to a parameter if there exists an instance of the problem with a fixed parameter such that solving this instance is harder than solving any problem in NP.
Theorem 6.1.
Elimination Distance to Dominated Clusters is para-NP-hard for parameter $k+\Delta$.
Proof 6.2.
Setting $k=0$ and $\Delta=3$ corresponds to Dominating Set on degree $3$ graphs. This problem is known to be NP-hard [23], which implies the statement.
Theorem 6.3.
Elimination Distance to Dominated Clusters is para-NP-hard for parameter $d$.
Proof 6.4.
Setting $d=0$ corresponds to TreeDepth. This problem is known to be NP-hard [32], which implies the statement.
W[2]-hardness for parameter $k+d$
A problem is W[2]-hard if a W[2]-hard problem reduces to it via an FPT-reduction.
Theorem 6.5.
Elimination Distance to Dominated Clusters is W[2]-hard for parameter $k+d$.
Proof 6.6.
The W[2]-hard Dominating Set [14] reduces to Elimination Distance to Dominated Clusters by simply setting $k=0$. This is an FPT-reduction.
No polynomial kernels for parameter $k+c+d$
A kernelization algorithm (or kernel) for a parameterized problem is a polynomial-time algorithm that maps each instance $(x,k)$ of a parameterized problem $L$ to an instance $(x^{\prime},k^{\prime})$ of $L$ such that
  • $(x,k)\in L\Leftrightarrow(x^{\prime},k^{\prime})\in L$, and
  • $|x^{\prime}|+k^{\prime}\leq f(k)$ for a computable function $f$.
A kernel is polynomial if $f$ is polynomial.
[…]
Recall that $c$ denotes the degeneracy of a graph. A graph $G$ is $c$-degenerate if every subgraph $H$ has a vertex of degree at most $c$ (in $H$).
Theorem 6.7.
TreeDepth is NP-hard on $2$-degenerate graphs.
Proof 6.8.
We reduce TreeDepth to TreeDepth on $2$-degenerate graphs. Given a graph $G$ and parameter $k$, we compute $H$ by replacing each edge of $G$ by two disjoint paths of length $2$ with the same endpoints. We set $k^{\prime}=k+1$ and prove that $G$ has treedepth at most $k$ if and only if $H$ has treedepth at most $k^{\prime}$. Observe that $H$ is $2$-degenerate.
Assume that $G$ has treedepth at most $k$ with an elimination tree $T$. We take the same elimination tree for $H$ and in the very last step eliminate the isolated subdivision vertices.
Vice versa assume that $H$ has treedepth at most $k+1$ with an elimination tree of depth at most $k+1$. Observe that every subdivision vertex must be comparable in the tree-order with both of its endpoints, as each edge must be controlled by the ancestor-descendant relation in the elimination tree. Now, if a sudivision vertex is eliminated before one of its endpoints, we can simply exchange the subdivision vertex with this endpoint and thereby obtain an elimination tree that is not deeper than the original one. In this way, we obtain an elimination tree where the subdivision vertices are on the lowest level. Then the tree with the subdivision vertices removed is an elimination tree for $G$ of depth at most $k$.
Corollary 6.9.
Elimination Distance to Dominated Clusters does not admit a polynomial kernel with respect to parameter $k$ even for fixed parameters $d=0$ and $c=2$.
7 Conclusion
We have studied the Dominated Cluster Deletion problem and resolved the open question by Bentert et al. [3] whether the problem is fixed-parameter tractable with respect to the parameters $k+d+c$, by proving uniform fixed-parameter tractability even for the smaller parameter $k+d+\ell$, where $\ell$ is the semi-ladder index of the input graph. We also introduced the more general Elimination Distance to Dominated Clusters problem and almost fully classified its parameterized and kernelization complexity. The most interesting missing part of the classification is whether the case $d=0$, which corresponds to TreeDepth, is NP-complete on graphs of bounded maximum degree. We conjecture that this is the case. Our proofs combine the main tools for addressing cut problems, the decomposition theorem into unbreakable parts by Cygan et al. [11] with the main tools for addressing domination type problems [20].
Appendix A Algorithm on bag graphs
This section is devoted to the proof of Proposition 5.3.
See 5.3
It closely follows from the proof of Theorem 4.15 in Section 4. The main differences are that we now work on an annotated version of the problem, and that the graph might not be unbreakable, only the set of interior vertices are (see Definition 5.2).
Definition A.1.
Let $G=(I,E)$ be a $(q,k,d)$-bag graph, $F$ a set with $E\subseteq F$, $R$ and $B$ additional colors on bag vertices. We call a set $S$ a skeleton for annotated dominated cluster deletion with parameters $k$ and $d$ if there is a superset $S^{\prime}\supseteq S$ of vertices satisfying:
  • $|S^{\prime}|\leq k$,
  • $S^{\prime}\subseteq G-F$,
  • every connected component of $G-S^{\prime}$ has a Red-Blue dominating set of size at most $d$, and
  • $S$ contains exactly the vertices of $S^{\prime}$ that have at least one neighbor in a component $C$ of $G-S$ where $|C\cap I|\leq q$, and $C$ contains at least one Red vertex.
We now define an operation that will be necessary to obtain a statement analogous to Lemma 4.6.
Definition A.2.
For a bag graph $G=(E,I)$, we say that two interior vertices $u,v$ are close if either $u$ and $v$ are adjacent, or if there is a path from $u$ to $v$ that only uses exterior vertices.
We then define the $q$-saturation $\mathsf{sat}_{q}(u)$ of an interior vertex $u$ as the set of all interior vertices that are close to $u$ if there are at most $q$ of them; or as the singleton $u$ otherwise. For a set $X$, $\mathsf{sat}_{q}(X):=\bigcup\limits_{u\in X}\mathsf{sat}_{q}(u)$.
For an integer $i$, we define the $q$-saturation at distance $i$ $\mathsf{sat}_{q}^{i}(X)$ of a vertex set $X$ as $\mathsf{sat}_{q}(X)$ if $i=1$, and $\mathsf{sat}_{q}(\mathsf{sat}_{q}^{i-1}(X))$. I.e. we apply the $q$-saturation operator $i$ times.
Intuitively from a bag vertex, we take all the bag vertices that are adjacent or connected through component vertices, but only if there are few of them (at most $q$).
Lemma A.3.
Let $q,k,d$ be integers, $G$ be a $(q,k)$-bag graph that is a positive instance to Annotated Dominated Cluster Deletion, then $G$ has a $3qd$ Red-Blue dominating set.
Proof A.4.
Let $(I,E)$ be the partition of $G$, and $S$ be a solution i.e. every connected component of $G-S$ has a $d$-Red-Blue dominating set. Let $C_{0}$ be the only connected component of $G-S$ that contains more than $q$ vertices of $I$. In $G-C_{0}$, there are at most $q$ interior vertices, which together with the $qd$ Red exterior vertices amount to at most $2qd$ Red vertices in $G-C_{0}$. So the Blue dominating set for $G-C_{0}$ is of size at most $2qd$ as well. Together with the $d$ Blue vertices dominating all Red vertices of $C_{0}$, we have a $3qd$ Red-Blue dominating set.
Lemma A.5 (analogous to Lemma 4.6).
Let $G=(I,E)$ be a $(q,k,d)$-bag graph with colors $F,R,B$, let $S$ be a skeleton for Annotated Dominated Cluster Deletion with parameters $k,d$, and let $X$ be a $3qd$ Red-Blue dominating set for $G$. Then there is a vertex of $v\in S$ that is in $sat_{q}^{q}(X)$.
Proof A.6.
If there is a vertex of $S$ in $X$, we are done as $X\subseteq\mathsf{sat}_{q}^{q}(X)$. Otherwise, let $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, let $v$ be a vertex of $S$ and $C$ the connected component of $G-S^{\prime}$ witnessing that $v\in S$ i.e. $|C\cap I|\leq q$, $C$ contains a Red vertex, and $v$ has a neighbor is $C$.
Let $u$ be a Blue vertex of $X$ that dominate a Red vertex of $C$. If $u$ is in $S$ then we are done. Otherwise, $u$ is in $C$ and there is a path in $C$ from $u$ to $v$. Each interior vertex $w$ of this path is in $C$ and every vertex close to $w$ is in $C$ or $S$, so there are at most $q$ close vertices and all are in $\mathsf{sat}_{q}(w)$. This concludes that $v\in\mathsf{sat}^{q}_{q}(u)$, hence $v\in\mathsf{sat}^{q}_{q}(X)$.
Definition A.7 (Skeleton algorithm for bag graph).
The colored skeleton-algorithm with parameters $q,k,d$ over a bag graph $G$ returns a family of sets of vertices, each set of size at most $k$.
First, it computes a dominating set $X$ of size at most $3qd$ (returning the empty family if there is no such set). Second, it computes the set $\mathsf{sat}^{q}_{q}(X)$. Third, it calls the algorithm with parameters $q,k-1,d$ on $G$ (and get a family $\mathcal{F}_{0}$), and on $G-v$ for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ (and get a family $\mathcal{F}_{v}$). And finally, for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ add $v$ to every set in $\mathcal{F}_{v}$, and return the union of all the families.
Remark A.8.
Note that for any interior vertex $v$, $G-v$ is a $(q,k-1,d)$-bag graph. Furthermore, if the input graph is a positive instance of Annotated Dominated Cluster Deletion, and if the algorithm guesses correctly a vertex $v$ of a solution, then $G-v$ is also a positive instance to Annotated Dominated Cluster Deletion with parameter $(k-1,d)$. So the algorithm can repeat until it guesses a skeleton $S$ of size at most $k^{\prime}$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
It remains to verify that $S$ can indeed be enlarged to a set $S^{\prime}$ of size at most $k$ with the property that every connected component of $G-S^{\prime}$ has a $d$-Red-Blue dominating set.
Note that by definition, there is at most one component $C_{0}$ of $G-S$ that contains more than $q$ interior vertices. So in $G-C_{0}$, there are at most $3qd+q$ many Red vertices ($3qd$ for the exterior vertices and $q$ for the interior ones). What remains to be done is hence the following.
  1. 1.
    Find an optimal solution for $G-S-C_{0}$. This can be done in time $f(k,d)$, because it contains at most $q$ interior vertices and only these can be deleted (so we can try all possibilities), and it contains at most $3qd+q$ Red vertices. As we can treat indifferently two Blue vertices with the same Red neighborhood, and since there are at most $2^{O(qd)}$ possible Red neighborhoods in time $f(q,k,d)$ we can go through all possibilities to dominate the Red vertices.
    Each such possibility yields a value $k^{\prime\prime}$ of vertices that need to be deleted. If $k^{\prime\prime}>k-k^{\prime}$ we reject the instance (in this branch). Remember that $k^{\prime}=|S|$.
  2. 2.
    Test whether an additional set $W\subseteq C_{0}-F$ of $k^{\prime\prime\prime}=k-k^{\prime}-k^{\prime\prime}$ vertices that can be deleted from $C_{0}$ such that $G[C_{0}-W]$ has a $d$-Red-Blue dominating set. Hence, the remaining problem is an instance of Partial Red-Blue Domination with parameter $k^{\prime\prime\prime},d$.
    If $G$ is a positive instance of Annotated Dominated Cluster Deletion, witnessed by some set $S^{\prime}$, and if $S$ is a skeleton of $S^{\prime}$, such a set $W$ exists (as $S^{\prime}\cap C_{0}$ is a candidate).
    If $t(\cdot)$ is a parameter satisfying 1, we can then test for the existence of such $W$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$. If no solution is found, we reject the instance (in this branch).
This concludes the proof of Proposition 5.3 and therefore Appendix A.
Appendix B Algorithm through unbreakable tree decomposition.
In this section, we assume that we have a fix graph $G$, and $(T,\mathsf{bag})$ a $(q,k)$-unbreakable tree decomposition of $G$. See Section 2 for the definition of an unbreakable tree decomposition, and Theorem 2.1 ([12, Theorem 10]) for the proof of its existence.
The goal of this section is to prove Theorem 1.1 using the fact that Annotated Dominated Cluster Deletion is FPT on bag graphs (Proposition 5.3).
See 1.1
B.1 Marks and unbreakable bag graph
Definition B.1.
Given a graph $G$ (with colors $F,R,B$), a subset of vertices $A$, and integers $k,d$, we define a $(k,d)$-mark of $A$ in $G$ as a tuple of the form:
  • disjoint subsets $S_{A},D_{A},U_{A}$ of $A$,
  • an integer $k_{A}\leq k$,
  • a partition ${\cal{P}}$ of $A-S_{A}$, and
  • for each element $P$ of the partition, an integer $p\leq d$.
A mark is said to be realized if there is a set $S\subseteq V(G)-F$ of at most $k$ vertices such that:
  • $S_{A}=S\cap A$,
  • $k_{A}=|S|-|S_{A}|$,
  • for every connected component $C$ of $G-S$ either :
    • $C\cap A=\emptyset$ and $C$ has a Red-Blue dominating set of size $d$, or
    • $C\cap A\neq\emptyset$, there is a $P\in{\cal{P}}$ such that $P=C\cap A$. Furthermore, there must exist a subset $D_{C}\subseteq C\cap B$ of size $p$ such that $D_{C}\cap P=D_{A}\cap P$, and $D_{C}$ dominates $R\cap(C-U_{A})$.
Intuitively, we mark how the set $A$ overlaps with the deleted vertices and the Blue dominating sets. The Blue dominating sets must dominate all Red vertices of $G-A$, and if some Red vertices of $A$ are not dominated, it is written so on $U_{A}$ (for un-dominated).
Definition B.2.
Given a graph $G$, a subset of vertices $A$, and integers $k,d$, the $(k,d)$-profile of $A$ in $G$ is the list of all realized $(k,d)$-marks of $A$ in $G$.
Lemma B.3.
The $(k,d)$-profile of a set $A$ of size $|A|=2^{\mathcal{O}(k)}$ in a graph $G$ has size at most $d^{2^{\mathcal{O}(k)}}$.
Proof B.4.
There are $\mathcal{O}(4^{|A|})$ possibilities to split the set $A$ into the disjoint subsets $S_{A},D_{A},U_{A}$, and the rest, and $k_{A}$ is an integer bounded by $k$. Furthermore, there are at most $|A|$ elements in the partition, each is given an integer bounded by $d$. This results in $\mathcal{O}(d^{|A|})$ possibilities.
Assuming that $|A|=2^{\mathcal{O}(k)}$, this yields profiles of size $d^{2^{\mathcal{O}(k)}}$.
The overall algorithm will inductively compute the profile of each cone of the tree decomposition (with respect to its adhesion). The profile of the root bag will contain the information to answer whether the graph is a positive instance to the Annotated Dominated Cluster Deletion problem. The main proof of this section is that we can compute the profile of a cone given the profile of every subtree. To achieve that we need to add some structure to the bags of the tree decomposition.
Definition B.5.
A $d$-gadget is the graph $H$ composed of $12d+$ 1 vertices $a,b_{1},\ldots,b_{d},b^{\prime}_{1}\ldots,b^{\prime}_{d}$, where for every $i\leq d$, $b_{i}$ is adjacent to $a$ and $b^{\prime}_{i}$.
A $d$-gadget is plugged to a set $A$ by making every vertex of $A$ adjacent to $a$.
Note that $H$ is the one-subdivided star with $d$ branch. With a slight abuse of notation, we also call $d$-gadget, any graph that are $d^{\prime}$-gadget with $d^{\prime}\leq d$.
Definition B.6.
Let $(T,\mathsf{bag})$ be a regular $(q,k)$-unbreakable tree decomposition of a graph $G$, $S\subseteq\mathsf{cone}(x)$ a set of at most $k$ vertices, and $d$ an integer. We note $\mathsf{bgraph}_{S}(x)$ of a node $x\in V(T)$ any graph that is obtained from $G[\mathsf{cone}(x)-S]$ by :
  • contracting each connected component $C$ of the set $\mathsf{comp}(y)-S$ that has a neighbor in $\mathsf{adh}(y)-S$ into a $d$-gadget $H_{y}^{C}$ and plugging $H_{y}^{C}$ to the vertices of $\mathsf{bag}(x)-S$ that are neighbors of $C$, and
  • adding at most $q$ many $d$-gadget plugged to a subset of the adhesion of $x$.
Lemma B.7.
Every $\mathsf{bgraph}_{S}(x)$ is a $(q,k-|S|,d)$-bag graph.
Proof B.8.
The vertices of $\mathsf{bgraph}_{S}(x)$ are split into $I$ the vertices of $\mathsf{bag}(x)$, and $E$ the union of all the $d$-gadgets. Clearly, every connected component of $G-I$ is a $d$-gadget and therefore has at most $12d+$ 1 vertices. Furthermore, this $d$-gadget is plugged to a subset of the adhesion of $x$ or the adhesion of a child of $x$, and therefore has at most $q$ neighbors in $I$.
For the last point, let $L,R$ be a separation of $\mathsf{bgraph}_{S}(x)$ of order at most $k-|S|$, then extend $L,R$ into $L^{\prime},R^{\prime}$ by first removing the $d$-gadget plugged to the adhesion of $x$, then adding $S$ to both $L$ and $R$, and finally replacing every $d$-gadget plugged to a subset of $\mathsf{adh}(y)$ by the component of $\mathsf{cone}(y)$ it represents.
We obtain that $(L^{\prime},R^{\prime})$ is a separation of order $k$ of $\mathsf{cone}(x)$, and therefore $|L^{\prime}\cap\mathsf{bag}(x)|\leq q$ or $|R^{\prime}\cap\mathsf{bag}(x)|\leq q$, hence with $I=\mathsf{bag}(x)$, either $|L\cap I|\leq q$, or $|R\cap I|\leq q$.
Note that $\mathsf{bgraph}_{S}(x)$ is not just one graph, but rather a family of graphs. Nonetheless, it is convenient to write $|\mathsf{bgraph}_{S}(x)|$ and $\|\mathsf{bgraph}_{S}(x)\|$ to denote the maximum amount of vertices (resp. of edges) that one of the $\mathsf{bgraph}_{S}(x)$ can have. We now provide bounds for these numbers.
Lemma B.9.
Let $(T,\mathsf{bag})$ be a regular $(q,k)$-unbreakable tree decomposition of a graph $G$. Then $\sum\limits_{x\in V(T)}|\mathsf{bgraph}(x)|\leq(4d+5)q|G|$ and $\sum\limits_{x\in V(T)}\|\mathsf{bgraph}(x)\|\leq\|G\|+((4d+5)q)^{2}|G|$.
Proof B.10.
For every $x\in V(T)$, $\mathsf{bag}(x)$ is the disjoint union of $\mathsf{adh}(x)$ and $\mathsf{mrg}(x)$. Every vertex of $G$ appears in exactly one margin of one node in the tree decomposition and each node of the tree decomposition has an adhesion of size at most $q$. Furthermore, each node is the child of at most one node and each connected component of a child can be divided in at most $q$ many connected components. Additionally, $d$-gadgets have size at most $12d+$ 1, and $|T|\leq|G|$, hence:
$\sum\limits_{x\in V(T)}|\mathsf{bgraph}(x)|\leq(2d+2)q|T|+|G|+(2d+2)q|T|\leq(4d+5)q|G|$
For the second bound, we consider the total number of vertices and edges in all bag graphs. Each edge in the margin of a node in the tree decomposition corresponds to exactly one edge in the graph $G$. Additionally, there are at most $q\cdot|\mathsf{bgraph}(x)|$ many edges in each bag graph that are incident to a vertex in the adhesion and there are at most $(2qd)\cdot|\mathsf{children}(x)|$ many edges in the $d$-gadgets plugged to the adhesion of the children of $x$ and $(2qd)$ edges in the $d$-gadgets plugged to the adhesion of $x$.
$\sum\limits_{x\in V(T)}\|\mathsf{bgraph}(x)\|$ $\leq\sum\limits_{x\in V(T)}|\mathsf{bgraph}(x)|+\|G\|+2dq\cdot\sum\limits_{x\in V(T)}(|\mathsf{bgraph}(x)|+|\mathsf{children}(x)|)$
$\leq(4d+5)q|G|+\|G\|+2dq((4d+5)q|G|+|T|)$
$\leq\|G\|+((4d+5)q)^{2}|G|$
This concludes the proof.
In the next lemma, we show that the ladder index of these bag graphs is bounded (assuming that it was already bounded in the original graph). This together with the fact that the semi-ladder index has 1 implies that the semi-ladder index also has 2.
Lemma B.11.
Let $\ell$ be an integer and $(T,\mathsf{bag})$ be regular $(q,k)$-unbreakable tree decomposition of a graph $G$ with semi-ladder index at most $\ell$. Then, for every node $x\in V(T)$ every bag graph $\mathsf{bgraph}_{S}(x)$ has semi-ladder index at most $q+\ell+4$.
Proof B.12.
Let $a_{1},\ldots,a_{t},b_{1},\ldots,b_{t}$ be a semi-ladder of $\mathsf{bgraph}_{S}(x)$. Note that in a $d$-gadget, only one vertex has degree greater than 2. Hence, the other vertices can only be among $a_{1},a_{2},a_{3}$ and $b_{t},b_{t-1},b_{t-2}$ and since the only high degree vertex of a $d$-gadget has at most $q$ neighbors that are in $\mathsf{bag}(x)$, we conclude that the vertices from the $d$-gadget can only be among $a_{1},\ldots,a_{q+4}$, and symmetrically in $b_{t-q-3},\ldots,b_{t}$. Therefore, $a_{q+5},\ldots a_{t},b_{1},\ldots,b_{t-q-4}$ is a semi-ladder of $G$. So $t-(q+5)+1\leq\ell$, hence $t\leq q+\ell+4$.
We state the next proposition with arbitrary parameter $t(\cdot)$ satisfying 2. By the previous lemma, the semi-ladder index $\ell$ can replace $t(\cdot)$.
Proposition B.13.
Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the tree decomposition and, for each child $y$ of $x$, the $(k,d)$-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 2. Or alternatively in time $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$.
Note that by Lemma B.9, the sum of the size of all bag graphs is linear in the size of the graph so that Proposition B.13 yields Theorem 5.4, and therefore Theorem 1.1.
B.2 Algorithm for Proposition B.13
Let $G$ be a graph, $(T,\mathsf{bag})$ a $(q,k)$-unbreakable tree decomposition and $x\in V(T)$ a node of this decomposition. We assume that we have access to the profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$ for every child $y$ of $x$. We fix a $(k,d)$-mark $M_{x}$ of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$ and test whether it is realized. In order to simplify a bit the notation, given a child $y$ of $x$ we will simply call a mark of $y$ a $(k,d)$-mark of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$.
[…]
Before going to the children, let $S_{x},D_{x},U_{x},k_{x},{\cal{P}}_{x},(p_{i})_{P_{i}\in{\cal{P}}_{x}}$ be the mark selected for $x$. We add $S_{x}$ to $S$, $k_{x}$ to $s$ and $D_{x}$ to $D$.
Branching through children’s marks. We now go over all children $y$ of $x$ and update our partial solution $W$. There are two ways our partial solution is updated, depending on the marks present in the profile of $y$. If the following mark is realized
$S_{y}=\mathsf{adh}(y)\cap S\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ D_{y}=\mathsf{adh}(y)\cap D\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ k_{y}=0\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \bigwedge\limits_{P_{i}\in{\cal{P}}_{y}}p_{i}=0,$
then we do not update our partial solution, this child is already taken care of, and we move to the next child of $x$.
Otherwise, our algorithm branches over all the realized marks of $y$. For each branch, considering a mark $M^{\prime}$, we check the consistency of $M^{\prime}$ with our partial solution, and if it is consistent we update our connected component labeling and our partial solution. The checks are :
  • $S\cap\mathsf{adh}(y)\subseteq S_{y}\subseteq G-F$, i.e. agrees on the vertices being already deleted and none are in $F$.
  • $D\cap\mathsf{adh}(y)\subseteq D_{y}\subseteq B$, i.e. agrees on the vertices already picked to be dominators, and they are Blue vertices.
  • $D\cap S_{y}=S\cap D_{y}=\emptyset$, i.e. a vertex cannot be both deleted and a dominator.
For the updates, first, add $A_{S}$ to $S$, and $k_{y}$ to $s$. Second, for every $P_{i}\in{{\cal{P}}}_{y}$ add $p_{i}$ to $c$.
If any of the previous checks failed, as well as if either $|S|+s>k$, or if $|D|+c>qd$, we terminate this branch of the algorithm.
Reducing to bag graphs. We have $S,s,D,c$ fixed and one selected mark for every child $y$ and one for $x$. We compute a bag graph $H$ as follows:
First, starts with $H:=G[\mathsf{bag}(x)-S]$, taking the edges and colors $F,R,B$ from $G$.
Second, for every child $y$ let $S_{y},D_{y},U_{y},k_{y},{\cal{P}}_{y}$ and an integer $p$ for each $P\in{\cal{P}}_{y}$ be the selected mark. Then:
  • remove $\mathsf{adh}(y)-U_{y}$ from $R$ (these vertices are dominated in $y$),
  • remove $N(D_{y})$ from $R$ (we mark these vertices as dominated), and
  • for every $P_{i}\in{\cal{P}}_{y}$ plug the $p_{i}$-gadget to the vertices of $P_{i}$.
Finally, let $S_{x},D_{x},U_{k},k_{x},{\cal{P}}_{x},$ and a $p$ for each $P\in{\cal{P}}_{x}$ be the selected mark for $\mathsf{adh}(x)$.
  • Remove $N(U_{x})$ from $B$ ($U_{x}$ shall not be dominated),
  • remove $N(D_{x})$ from $R$ (we mark these vertices as dominated), and
  • for every $P\in{\cal{P}}_{x}$ plug the $(d-p)$-gadget to the vertices of $P$.
We are now done building a bag graph $H$, of the form $\mathsf{bgraph}_{S}(x)$. We now conclude by testing whether $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion with Proposition 5.3.
Claim 1 (Correctness of the algorithm).
The mark $M_{x}$ of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$ is realized if and only if $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion.
Proof B.14.
For the first direction, assume that $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion, then the algorithm of Proposition 5.3 produced a set $S^{\prime}$ of size $k_{x}-|S|-s$, which together with $S$ and the sets $S_{y}$ (of cumulative size $s$) selected for the children satisfies that every connected component of $\mathsf{cone}(x)-S-S^{\prime}-\bigcup\limits_{y\in\mathsf{children}(x)}S_{y}$ that does not intersect $\mathsf{adh}(x)$ has a Red-Blue dominating set of size $d$. For the component intersecting $\mathsf{adh}(x)$ the Red-Blue dominating set of size $d$ must contain the $d-p$ Blue vertices of the $(d-p)$-gadget; hence the component has a Red-Blue dominating set of size $p$.
For the second direction, assume that the mark is realized, hence there is a set $\hat{S}$ on $\mathsf{cone}(x)$ witnessing it this mark is realized, together with a set of Blue vertices $D_{C}$ for every connected component of $\mathsf{cone}(x)-\hat{S}$. When branching over the marks of the children of $x$, the algorithm select the mark for which $\hat{S}\cap\mathsf{cone}(y)$ is a witness. The built graph $H$ is then a positive instance witnessed by $\hat{S}\cap\mathsf{bag}(x)$.
B.3 Running time and Black-White trees
In this section, we explore the tree produced by our backtracking algorithm. We will show that this tree must have a specific structure that we call Black-White (with additional parameters), and such trees must be of linear size and have a constant number of leaves.
Definition B.15.
A tree is called Black-White if every node of the tree is colored either Black or White, and every non-leaf node either has exactly one White child or only Black children.
The degree of a Black-White tree is the maximum number of Black children a node has in the tree. The budget of a Black-White tree is the maximum number of Black vertices in a root-to-leaf path.
Lemma B.16.
Any Black-White tree of degree $\alpha$, budget $\beta$, and depth $n$ has size at most $\alpha^{\beta}\cdot n$ and the number of leaves is at most $\alpha^{\beta}$.
Proof B.17.
The bound on the number of leafs is straightforward. For the size, notice that whenever a White vertex has some Black children, it only increases the size of the tree to switch them (i.e. the parent of the White vertex now has several Black children, each with one White child), and this operation maintains the degree, budget, and depth of the tree. Therefore, the maximal Black-White tree is a complete Black tree of degree $\alpha$ and of depth $\beta$, where each leaf is attached to a White path of length $n-\beta$, for which the size is at most $\alpha^{\beta}\cdot n$.
When doing the backtracking, it is key that if a child is satisfied with the given partial solution we do not try to branch and make an impactful decision. Otherwise, the tree of partial solutions grows too much and is not linear. In the context of Black-White tree, that would be like allowing one White child and some Black children.
Lemma B.18.
The branching tree of backtracking algorithm of Section B.2 can be labelled to be a Black-White tree of degree $d^{2^{\mathcal{O}(k)}}$ and of budget $(k+qd)$.
Proof B.19.
When we branch over the children we only have $d^{2^{\mathcal{O}}(k)}$ many marks to consider (see Lemma B.3). This bounds the degree of our Black-White tree.
Furthermore, every time that we branch, this increases at least one of $|S|$, $s$, $|D|$, or $c$. As we terminate the algorithm as soon as $|S|+s>k$ or $|D|+c>qd$, the budget remains bounded by $k+qd$.
Going back to the algorithm for Proposition B.13 presented in Section B.2, we just showed that the number of possible branches to consider is only a function of $q,k$, and $d$. And since each branch then perform a single call to Proposition 5.3, this concludes the proof of Theorem 5.4.
Appendix C Annotated version for Elimination Distance to Dominated Clusters
Similarly to Section 5, we extend Theorem 1.2 to an annotated version.
Definition C.1 (Annotated Elimination Distance to Dominated Clusters).
Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Elimination Distance to Dominated Clusters problem is to find a tree-structured set $S\subseteq G-F$ of elimination depth $k$ such that every connected component $C$ of $G-S$ has a Red-Blue dominating set of size at most $d$ (i.e. $d$ Blue vertices dominating all Red vertices).
Definition C.2.
Given a graph $G$, a subset of vertices $A$, and integers $k,d$, we define a $(k,d)$-extended-mark of $A$ in $G$ as a tuple of the form:
  • disjoint subsets $L_{A}^{1},\ldots,L_{A}^{k},D_{A},U_{A}$ of $A$,
  • a partition $\cal P$ of $A-L_{A}$, where $L_{A}=L_{A}^{1}\cup\ldots\cup L_{A}^{k}$, and
  • for each element $P$ of the partition:
    • an integer $p\leq d$,
    • booleans $r^{1},\ldots,r^{k}$.
An extended-mark is said to be realized if there is a set $S$, partitioned into sets $L^{1},\ldots,L^{k}$ of vertices such that:
  • for every $i\leq k$ and vertices $u,v$ in $L^{i}$ every $u-v$ path must use a vertex in $L^{j<i}$,
  • $L_{A}^{i}=L^{i}\cap A$, and
  • for every connected component $C$ of $G-S$ either:
    • $C\cap A=\emptyset$ and $C$ has a Red-Blue dominating set of size $d$, or
    • $C\cap A\neq\emptyset$ and there is a $P\in\cal P$ such that $P=C\cap A$. Furthermore, there must exist a subset $D_{C}\subseteq C\cap B$ of size $p$ such that $D_{C}\cap P=D_{A}\cap P$ and $D_{C}$ dominates $R\cap(C-U_{A})$.
      And $r^{i}=True$ if and only if there is a path from any vertex of $P$ to any vertex of $L^{i}$ that is not using vertices of $L^{j<i}$.
The first point corresponds to the fact that $S$ is tree-structured with elimination depth $k$. The sets $L^{i}_{A}$ correspond to the layer of the structure.
Definition C.3.
Given a graph $G$, a subset of vertices $A$, and integers $k,d$, the $(k,d)$-extended-profile of $A$ in $G$ is the list of all realized $(k,d)$-extended-marks of $A$ in $G$.
The proof of the next observation is similar to the one of Lemma B.3. {observation} The $(k,d)$-extended-profile of a set $A$ of size at most $q$ in a graph $G$ has size at most $(d+k)^{2^{\mathcal{O}(k)}}$.
We are now going to work with the same definition of gadget and bag graphs (see Definitions B.5 and B.6).
Proposition C.4.
Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the decomposition and, for each child $y$ of $x$, the $(k,d)$-extended-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-extended-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time either $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$, or $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t(\cdot)$ is a graph parameter satisfying 2.
Equivalently, the running time is $f(q,k,d,t_{1}(\mathsf{bgraph}(x)))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t_{1}(\cdot)$ is a graph parameter satisfying 1. Note that by Lemma B.9, the sum of the size of all bag graphs is linear in the size of the graph so Proposition C.4 yields our main result, Theorem 1.2.
C.1 Algorithm for Proposition C.4
Let $G$ be a graph, $(T,\mathsf{bag})$ a $(q,k)$-unbreakable tree decomposition and $x\in V(T)$ a node of this decomposition. We assume that we have access to the extended-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$ for every child $y$ of $x$. We fix a $(k,d)$-extended-mark $M_{x}$ of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$ and test whether it is realized. In order to simplify a bit the notation, given a child $y$ of $x$ we will simply call an emark of $y$ a $(k,d)$-extended-mark of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$.
Initialization
We now design a backtracking algorithm that start with a partial solution $W=\{S^{1}=\ldots=S^{k}=\emptyset,$ $D=\emptyset,s^{1}=\ldots=s^{k}=0,c=0\}$. Intuitively, $S^{i}$ (resp. $D$) contains the vertices of $\mathsf{bag}(x)$ in the $i$th layer of the tree structure (resp. picked as dominators), and $s^{i}$ counts the vertices to be deleted in the descendant of $x$ that are in layer $L^{i}$ and reachable by a path not using $L^{j<i}$. Finally, the role of $c$ is to count the dominators from the children of $x$ that are in a connected component overlapping both $x$ and $y$.
Before going to the children, let $L^{1}_{x},\ldots,L^{k}_{x},D_{x},U_{x},{\cal{P}}_{x},(p_{i})_{P_{i}\in{\cal{P}}_{x}},r^{1}_{x},\ldots,r^{k}_{x}$ be the mark selected for $x$. We add $L^{i}_{x}$ to $S$, and $D_{x}$ to $D$.
Children
We now go over all children $y$ of $x$ and update our partial solution $W$. There are two ways our partial solution is updated, depending on the extended-marks present in the extended-profile of $y$. If the following extended-mark is realized
${L^{i}_{y}}=\mathsf{adh}(y)\cap S^{i}\leavevmode\nobreak\ \leavevmode\nobreak\ (\text{for all }i\leq k);\leavevmode\nobreak\ \leavevmode\nobreak\ D_{y}=\mathsf{adh}(y)\cap D\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak$
$r^{i}=False\leavevmode\nobreak\ \leavevmode\nobreak\ (\text{for all }i\leq k);\leavevmode\nobreak\ \leavevmode\nobreak\ \bigwedge\limits_{P_{j}\in{\cal{P}}_{y}}p_{j}=0,$
then we do not update our partial solution, this child is already taken care of, and we move to the next child of $x$.
Otherwise, our algorithm branches over all the realized extended-marks of $y$. For each branch, considering an extended-mark $M_{y}$, we check the consistency of $M_{y}$ with our partial solution, and if it is consistent we update our connected component labeling and our partial solution. The checks are :
  • $S^{i}\cap\mathsf{adh}(y)\subseteq{L^{i}_{y}}$, i.e. agrees on the vertices being already deleted, and their layer.
  • $D\cap\mathsf{adh}(y)\subseteq D_{y}$, i.e. agrees on the vertices already picked to be dominators.
  • $S^{i}\cap D_{y}=D\cap{L^{i}_{y}}=S^{i}\cap{L_{y}^{j\neq i}}=\emptyset$, i.e. a vertex cannot be both deleted and a dominator and if deleted only on a single layer.
As for the updates, first, add $L_{y}^{i}$ to $S^{i}$, and increment $s^{i}$ if $r^{i}=True$. Second, for every $P_{i}\in{{\cal{P}}_{y}}$ add $p_{i}$ to $c$.
If any of the previous checks failed, as well as if either $|S^{i}|+s^{i}>qk$ (for any $i\leq k$), or if $|D|+c>qd$, we terminate this branch of the algorithm.
Bag graphs and skeletons
We now have a fixed set $S=\bigcup\limits_{i\leq k}S^{i}$, and similarly to Section B.2, we compute the bag graph $H_{x}=\mathsf{bgraph}_{S}(x)$. Similarly to Appendix A (see Lemma A.3) positive instances have a small Red-Blue dominating set. This transfers to the elimination distance variation, as it was the case in Lemma 4.21 for unbreakable graphs. So the colored version of skeleton-algorithm (see Definition A.7) can also be used here.
Then we therefore compute every skeleton $S^{\prime}$ with parameter $k$ and $3qd$ of $H_{x}$. We then guess in which $S^{i}$ we assign each element of $S$ (which modifies $S$). We terminate this branch of the algorithm if $|S^{i}|+s_{i}>qk$, or if $S^{\prime}$ contains a vertex that lies in the adhesion of a child of $x$ (meaning that the selected mark for this child is inconsistent with the selected skeleton).
This step can be performed in time $f(q,k,d,t_{1}(H_{x}))\cdot|H_{x}|$, where $t_{1}(\cdot)$ satisfies 1, meaning in time $f(q,k,d,t_{2}(G))\cdot|H_{x}|$, where $t_{2}(\cdot)$ satisfies 2. Alternatively, in time $f(q,k,d)\cdot|H_{x}|^{\mathcal{O}(d)}$, in which case we also store the set $D^{\prime}$ that dominates the large connected component. See Lemma 4.25, which can be adapted to bag graphs.
Fixing the small connected components
We now define $C_{0}$ as the only connected component of $H_{x}-S$ that contains more than $q$ vertices of $\mathsf{bag}(x)$. We branch over every choice $S^{\prime\prime}\subseteq\mathsf{bag}(x)-(S^{\prime}\cup C_{0})$ such that $S^{\prime\prime}\cup S$ still has a tree structure of elimination depth $k$. This creates connected components $C_{1},\ldots,C_{q}$ of $\mathsf{bag}(x)-(S\cup S^{\prime\prime}\cup C_{0})$. We guess for each vertex of $S^{\prime\prime}$ in which $S_{i}$ it is assigned.
For every $i>0$, we define $D_{i}:=D\cap C_{i}\cup D^{\prime}_{i}$, where $D$ is the set computed at the end of the Children step and $D^{\prime}_{i}$ is a guessed subset of at most $d$ Blue vertices of $C_{i}$ (we branch over all possibilities).
Partial domination in the large connected component
We now compute a partial dominating set in the large connected component. This is achieved as for Corollary 3.4 for which the colors $F,R,B$ are defined as:
  • $F$: Starts with $F$ from $G$, add every $d$-gadgets, and the adhesion of every $y$, children of $x$, together with the adhesion of $x$.
  • $R$: Remove from $R$ every vertex already dominated by the current value of $D$.
  • $B$: Remove from $B$ the vertices of $\mathsf{adh}(x)$ that are in $U_{x}$ and their neighborhood (as these vertices shall not be dominated).
We check the existence of a solution for every pair of parameters $k^{\prime}\leq k$, and $d^{\prime}\leq d$ (we branch over the choice of $k^{\prime}$ and $d^{\prime}$).
Alternatively, we add $D^{\prime}$ to $D_{0}$ and check that it dominates $C_{0}$. In the case $D^{\prime}$ was computed in time $f(q,k,d)\cdot|H_{x}|^{\mathcal{O}(d)}$.
Sanity checks
We no longer branch and only check whether this branch satisfies our requirements.
First check that in every small connected component $C_{i>0}$, every Red vertex is dominated by $D_{i}$. We then check that the connected components of $\mathsf{cone}(x)$ all have small Red-Blue domination sets. I.e. there could be connected components $C_{i}$, $C_{j}$ of $\mathsf{bag}(x)$ that both have a neighbor in a $d$-gadget in $H_{x}$. We call two such connected components related.
We can then define the related graph on the connected components by going through all the $d$-gadget. This yields a partition of the connected components where each family is a maximal subset of related connected components.
Then for every $I\subset\{0,\ldots q\}$ of this partition we add up the size of $|D_{i\in I}|$, as well as the relevant $p_{j}$ from the selected marks of the children of $x$ that are adjacent to $C_{i\in I}$.
After that, we check the consistency of the $S^{1},\ldots,S^{k}$ i.e. no two vertices of $S^{1}$ are connected in the graphs, then no two vertices of $S^{2}$ can be connected not using $S^{1}$ etc. This can be done in linear time: Simply brute force the paths in the small components, and, in the big component, every two vertices are connected not using $S$ (otherwise the large component is split).
We finally check that the current solution matches the selected extended-mark $M_{x}$ that we are trying to realize. So we check that the number of deleted vertices in each of the $k$ layer correspond to the values in $r^{i}_{x}$ (for each $i\leq k$), and similarly for the number of picked dominators per family of connected components $C_{i\in I}$ for each relevant $I$, and check that the partition ${\cal{P}}_{x}$ agrees on this split.
If all the checks were successful, the algorithm claims that the mark is realized.
C.2 Correctness and running time
Both are straightforward adaptation of what was done for Annotated Dominated Cluster Deletion. For example Section B.3 also exactly explain the running time of the current algorithm.
For correctness, pick a realized $(k,d)$-extended-mark $M_{x}$ of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$. Let $\hat{S}$ be the witness set (together with its tree structure). And for every connected component $C$ of $\mathsf{cone}(x)-S$ a set $D$ whose existence is guarantied as the mark is realized.
When going through the ’Children’ steps, there is a branch of the algorithm that always took the extended-mark $M_{y}$ of $y$ that is realized and for which $S$ and the witness. Similarly, there is a choice of a skeleton that is exactly the skeleton of $S$. The other vertices of $S^{\prime}$ are either guessed from the small connected component, or with the partial domination in $C_{0}$. So the algorithm agrees.
The other direction is straightforward as the algorithm produced the witness sets and checks the existence of its tree structure.

### v2

@@
 Abstract
-In the Dominated Cluster Deletion problem we are given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether there exists a set of at most $k$ vertices whose removal results in a graph in which each connected component has a dominating set of size at most $d$.
-In the Elimination Distance to Dominated Clusters problem we are again given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether we can recursively delete vertices up to depth $k$ such that each remaining connected component has a dominating set of size at most $d$.
+In the Dominated Cluster Deletion problem, we are given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether there exists a set of at most $k$ vertices whose removal results in a graph in which each connected component has a dominating set of size at most $d$.
+In the Elimination Distance to Dominated Clusters problem, we are again given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether we can recursively delete vertices up to depth $k$ such that each remaining connected component has a dominating set of size at most $d$.
 Bentert et al.
@@
 We furthermore show that both problems are FPT when parameterized by $k+d+\ell$, where $\ell$ is the semi-ladder index of the input graph, a parameter that is upper bounded and may be much smaller than the degeneracy $c$, positively answering the open question of Bentert et al.
-We almost complete the picture by providing an almost full classification for the parameterized complexity and kernelization complexity of Elimination Distance to Dominated Clusters.
+We further complete the picture by providing an almost full classification for the parameterized complexity and kernelization complexity of Elimination Distance to Dominated Clusters.
 The one difficult base case that remains open is whether Treedepth (the case $d=0$) is NP-hard on graphs of bounded maximum degree.
@@
 1 Introduction
-[…]
+Assume that $\mathcal{Q}$ is a graph problem that is hard to solve on general graphs but $\mathscr{C}$ is a class of graphs where $\mathcal{Q}$ can be solved efficiently.
+In many cases, we can also solve $\mathcal{Q}$ efficiently on instances that are close to the instances of $\mathscr{C}$, say, on graphs that belong to $\mathscr{C}$ after the deletion of a small number of vertices.
+Guo et al.
+[25] formalized this concept under the name distance from triviality.
+For example, the size of a minimum vertex cover is the distance to the class of edgeless graphs and the size of a minimum feedback vertex set is the distance to the class of acyclic graphs.
+Cluster Vertex Deletion is the problem to determine for a given graph and number $k$ whether there exists a set of at most $k$ vertices whose deletion results in a cluster graph, that is, a graph in which every connected component is a clique [5, 34].
+Many generalizations and special cases of Cluster Vertex Deletion have been studied.
+For example, one may require that each connected component in the resulting graph is an $s$-club (a graph of diameter at most s) [21, 28] or an $s$-plex (a graph in which each vertex has degree at least nsn-s) [26, 35].
+Note that the special case where each clique in the solution graph has to be of size one is again the Vertex Cover problem.
+Very recently, Bentert et al.
+[4] studied the following Dominated Cluster Deletion problem: Given a graph $G$ and integer parameters $k$ and $d$, does there exist a set $S$ of at most $k$ vertices such that each connected component of GSG-S can be dominated by at most $d$ vertices?
+Bentert et al.
+provided an almost complete classification of the parameterized complexity and kernelization complexity of Dominated Cluster Deletion with respect to the parameters $k$, $d$, $c$, and $\Delta$, where $c$ and $\Delta$ are the degeneracy and the maximum degree of the input graph, respectively.
+They proved the following results: Dominated Cluster Deletion is para-NP-hard for parameters $k+\Delta$ and $d+\Delta$, $W[2]$-hard for parameter $k+d$ and admits a non-uniform algorithm with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$, and is non-uniformly fixed-parameter tractable with respect to parameter $k+d+\Delta$.
+They left as an open problem whether the problem is fixed-parameter tractable with respect to the parameter $k+d+c$.
+They also showed that the problem does not admit a polynomial kernel even for $d=1$ with respect to the parameter $k+c$, or with respect to the parameter $k+d+\Delta$, unless $\textsf{NP}\subseteq\textsf{coNP/poly}$.
 A related measure is the elimination distance to a class $\mathscr{C}$, which measures the number of recursive deletions of vertices needed for a graph $G$ to become a member of $\mathscr{C}$.
-Formally, a graph $G$ has elimination distance 00 to $\mathscr{C}$ if $G\in\mathscr{C}$, and otherwise elimination distance $k+1$, if in every connected component of $G$, we can delete a vertex such that the resulting graph has elimination distance $k$ to $\mathscr{C}$.
+Formally, a graph $G$ has elimination distance 0 to $\mathscr{C}$ if $G\in\mathscr{C}$, and otherwise elimination distance $k+1$, if in every connected component of $G$, we can delete a vertex such that the resulting graph has elimination distance $k$ to $\mathscr{C}$.
 Note that elimination distance to the class containing only the empty graph corresponds to the well-known measure treedepth.
@@
 The related concept of recursive backdoors has recently been studied in the context of efficient SAT solving [18, 30].
+To the best of our knowledge, we are the first to study elimination distance to classes with small dominating sets.
 In this work, we answer the open questions of Bentert et al.
@@
 positively.
-Our main contribution are the uniform FPT algorithms for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on semi-ladder-free graphs.
+Our main contributions are the uniform FPT algorithms for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on semi-ladder-free graphs.
 Theorem 1.1.
@@
   6. 6.
-[…]
+    The problem does not admit a polynomial kernel with respect to parameter $k$ even for fixed parameters $d=0$ and $c=2$ unless $\textsf{NP}\subseteq\textsf{coNP/poly}$; The parameter $d=0$ corresponds to Treedepth.
+We prove the simple observation that Treedepth is NP-complete on $2$-degenerate graphs.
+It is then trivial using an AND-cross-composition to prove that Treedepth on $2$-degenerate graphs does not admit a polynomial kernel unless $\textsf{NP}\subseteq\textsf{coNP/poly}$.
 The one case that we could not resolve is the parameter $d+\Delta$.
@@
 [29] proved a powerful meta theorem for logically defined graph properties.
-Whenever a CMSO-property (monadic second-order logic with modulo counting predicates) can be solved efficiently on unbreakable graphs (for appropriately chosen parameters $k$ and $q$), then it can also be solved efficiently on general graphs.
+Whenever a CMSO-property (monadic second-order logic with modulo counting predicates) can be solved efficiently on unbreakable graphs (for appropriately chosen parameters $k$ and $q$), then it can also be decided efficiently on general graphs.
 There is a small caveat though.
@@
 rely on this framework and show how to solve Dominated Cluster Deletion on unbreakable graphs in time $f(k,d)\cdot n^{\mathcal{O}(d)}$.
-The key combinatorial observation is that in $(q,k)$-unbreakable graphs, by the deletion of at most $k$ vertices leads to one giant component and a bounded number of remaining small components.
+The key combinatorial observation is that in $(q,k)$-unbreakable graphs, the deletion of at most $k$ vertices leads to one giant component and a bounded number of remaining small components.
 One can then guess a dominating set for the giant component in time $\mathcal{O}(n^{d})$ and argue combinatorially to efficiently find the vertices to be deleted.
@@
 one obtains the non-uniform algorithm for general graphs with the desired running time.
-We remark that Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are in fact a first-order definable problems, that is, for all parameters $k,d$ there exist first-order formulas of length $\mathcal{O}(k,d)$ defining the problems.
+We remark that Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are in fact first-order definable problems, that is, for all parameters $k,d$, there exist first-order formulas of length $\mathcal{O}_{k,d}(1)$ defining the problems.
 Hence, the problems are fixed-parameter tractable on all classes on which first-order model checking is fixed-parameter tractable, e.g.
 on nowhere dense classes [24] and even monadically stable graph classes [16, 17], as well as on classes with bounded cliquewidth [11] (and classes with bounded twinwidth [9] when contraction sequences are given with the input).
-Note that first-order logic cannot define connected components in general, however, the connected components arising in the recursive deletion of elements must have small diameter when they can be dominated by few vertices, hence, connected components in positive instances of Dominated Cluster Deletion and Elimination Distance to Dominated Clusters become first-order definable.
+Note that first-order logic cannot define connected components in general.
+However, the connected components arising in the recursive deletion of elements must have small diameter when they can be dominated by few vertices.
+Hence, connected components in positive instances of Dominated Cluster Deletion and Elimination Distance to Dominated Clusters become first-order definable.
 Thus, on these classes one obtains algorithms with running time $f(k,d)\cdot n^{\mathcal{O}(1)}$ by the meta theorems for first-order logic.
-It is conjectured that monadically dependent classes are the most general hereditary graph classes that admit efficient first-order model checking, we hence included them in the inclusion diagram below.
+It is conjectured that monadically dependent classes are the most general hereditary graph classes that admit efficient first-order model checking, we hence included them in the inclusion diagram, see Figure˜1.
+Figure 1: Relation between the discussed graph classes.
+Arrows indicate inclusion.
+Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are fixed-parameter tractable on monadically stable classes by a meta theorem for first-order logic.
+The same is true for bounded twinwidth classes when contraction sequences are given with the input.
+Our main result is the fixed-parameter tractability on semi-ladder-free graphs.
 This leaves two main questions.
@@
 [4] to reduce to unbreakable graphs.
-We first observe that on unbreakable graphs, the notion of elimination distance almost collapses to distance to triviality, which allows us to treat Dominated Cluster Deletion and Elimination Distance to Dominated Clusters in a similar way.
-In order to avoid the non-uniform approach of Lokshtanov et al., we use a decomposition theorem due to Cygan et al.
+We first observe that on unbreakable graphs, the notion of elimination distance almost collapses to distance to triviality, enabling us to treat Dominated Cluster Deletion and Elimination Distance to Dominated Clusters in a similar way.
+This idea has already been examined for hereditary classes in several papers see e.g.
+[1].
+Note that being a dominated cluster is not a hereditary notion.
+In order to avoid the non-uniform approach of Lokshtanov et al.
+[29], we use a decomposition theorem due to Cygan et al.
 [13].
@@
 We are given a graph with three sets $F$, $R$ and $B$ and parameters $d$ and $k$.
-We ask whether there exist $k$ vertices of $G-F$ ($F$ stands for forbidden) that may be deleted such that the remaining graph has a Red-Blue dominating set of size at most $d$, i.e., the vertices of $R$ can be dominated by $d$ vertices of $B$ ($R$ and $B$ stand for Red and Blue).
+We ask whether there exist $k$ vertices of GFG-F ($F$ stands for forbidden) that may be deleted such that the remaining graph has a Red-Blue dominating set of size at most $d$, i.e.
+the vertices of $R$ can be dominated by $d$ vertices of $B$ ($R$ and $B$ stand for Red and Blue).
 Formally, the problem is defined as follows.
-Given a graph $G$ with unary predicates $F\subseteq V(G)$, $R\subseteq V(G)$, and $B\subseteq V(G)$ and parameters $d$ and $k$, does there exist a set $D\subseteq B$ of size at most $d$ and a set $X\subseteq V(G)-F$ of size at most $k$ such that $G[R-X]$ is dominated by $D$?
+Given a graph $G$ with (not necessarily disjoint) sets $F\subseteq V(G)$, $R\subseteq V(G)$, and $B\subseteq V(G)$ and parameters $d$ and $k$, does there exist a set $X\subseteq V(G)-F$ of size at most $k$ and a set $D\subseteq B-X$ of size at most $d$ such that G[RX]G[R-X] is dominated by $D$?
 The problem can obviously be solved in time $n^{\mathcal{O}(d)}$ (guess the dominating set and check if at most $k$ Red vertices remain undominated).
-By our dynamic programming approach we hence get a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ in general graphs.
+By our dynamic programming approach, we hence get a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ in general graphs.
 We then turn our attention to restricted graph classes where (Annotated) Partial Domination can be solved more efficiently.
 Using an approach of Fabiański et al.
-[19] we observe that (Annotated) Partial Domination can be solved efficiently on all semi-ladder-free graph classes, which yields our second main result.
-Figure 1: Relation between the discussed graph classes.
-Arrows indicate inclusion.
-Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are fixed-parameter tractable on monadically stable classes by a meta theorem for first-order logic.
-The same is true for bounded twinwidth classes when contraction sequences are given with the input.
-Our main result is the fixed-parameter tractability on semi-ladder-free graphs.
+[20], we observe that (Annotated) Partial Domination can be solved efficiently on all semi-ladder-free graph classes, which yields our second main result.
 Semi-ladder-free are very general graph classes.
@@
 On the other hand, the class of all $1$-subdivided cliques is semi-ladder-free (and even $2$-degenerate) but is neither monadically stable nor has bounded twinwidth.
-As mentioned, we require that (Annotated) Partial Domination needs to be solved on a slightly modified graph class (on the bag graphs), however, the property of being semi-ladder-free is preserved by this modification, so that ultimately we conclude our main results, Theorem 1.1 and Theorem 1.2.
+As mentioned, we require that (Annotated) Partial Domination needs to be solved on a slightly modified graph class (on the bag graphs), however, the property of being semi-ladder-free is preserved by this modification, so that ultimately we conclude our main results, Theorem˜1.1 and Theorem˜1.2.
 Organization.
 The paper is organized as follows.
-After recalling the necessary notation in Section 2, we talk more precisely about partial domination and the graph parameters enabling FPT evaluation in Section 3.
-We then present how to solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graph with bounded semi-ladder index in Section 4.
-After that, in Section 5, we sketch the uniform algorithm on general semi-ladder-free graphs for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$.
-The details are in appendix.
-The hardness results conclude the paper in Section 6.
+After recalling the necessary notation in Section˜2, we talk more precisely about partial domination and the graph parameters enabling FPT evaluation in Section˜3.
+We then present how to solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graph with bounded semi-ladder index in Section˜4.
+After that, in Section˜5, we sketch the uniform algorithm on general semi-ladder-free graphs for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$.
+The details are in the appendix.
+The hardness results conclude the paper in Section˜6.
 2 Preliminaries
@@
 A connected component of $G$ is a maximal connected subgraph of $G$.
-For a vertex subset $X\subseteq V(G)$ we write $G[X]$ for the subgraph of $G$ induced by $X$.
-We write $G-X$ for $G[V(G)\setminus X]$ and for singleton sets $\{v\}$ we write $G-v$ instead of $G-\{v\}$.
+For a vertex subset $X\subseteq V(G)$, we write $G[X]$ for the subgraph of $G$ induced by $X$.
+We write GXG-X for $G[V(G)\setminus X]$ and for singleton sets $\{v\}$ we write GvG-v instead of $G-\{v\}$.
 We denote the open neighborhood of a vertex $v$ by $N(v)$ and the closed neighborhood (including the vertex $v$) by $N[v]$.
-For a set $Y\subseteq V(G)$ we write $N[Y]$ for $\bigcup_{v\in Y}N[y]$.
+For a set $Y\subseteq V(G)$, we write $N[Y]$ for $\bigcup_{v\in Y}N[y]$.
 Semi-ladders.
@@
 The semi-ladder index of a graph is the maximum order of a semi-ladder that it contains.
-A class of graphs has bounded semi-ladder index or is called semi-ladder-free if there is some $\ell$ such that the semi-ladder index of all of its members is bounded by $\ell$, see Figure 2.
+A class of graphs has bounded semi-ladder index or is called semi-ladder-free if there is some $\ell$ such that the semi-ladder index of all of its members is bounded by $\ell$, see Figure˜2.
 $a_{1}$$a_{2}$$a_{3}$$a_{4}$$b_{1}$$b_{2}$$b_{3}$$b_{4}$
@@
 There exists an integer $q\in 2^{\mathcal{O}(k^{2})}$ and an algorithm that computes in time $q\cdot|G|^{2}\|G\|$ a tree decomposition $(T,\mathsf{bag})$ of $G$ with at most $|G|$ nodes such that the following conditions hold:
-  • For every node $x\in V(T)$, the bag $\mathsf{bag}(x)$ is $(q,k)$-unbreakable in the subgraph induced by $G[\mathsf{cone}(x)]$.
+  • For every node $x\in V(T)$, the bag $\mathsf{bag}(x)$ is $(q,k)$-unbreakable in the subgraph $G[\mathsf{cone}(x)]$.
   • For every node $x\in V(T)$, the adhesion $\mathsf{adh}(x)$ is of size at most $q$.
@@
   • every vertex of $\mathsf{adh}(x)$ has a neighbor in $\mathsf{comp}(x)$.
-We may further assume that the tree decomposition we get from Theorem 2.1 is regular, see the construction of Lemma 2.8 of [7] and the discussion in [31].
+We may further assume that the tree decomposition we get from Theorem˜2.1 is regular, see the construction of Lemma 2.8 of [8] and the discussion in [31].
 Parameterized complexity.
@@
 3 Partial domination and FPT evaluation
-We show how to efficiently solve Annotated Partial Domination problem on semi-ladder-free graphs.
-This problem falls into the framework of domination type problems on semi-ladder-free graphs introduced by Fabiański et al.
+We show how to efficiently solve the Annotated Partial Domination problem on semi-ladder-free graphs.
+This problem falls into the framework of domination-type problems on semi-ladder-free graphs introduced by Fabiański et al.
 [20].
@@
 A domination-type problem is a formula $\exists\bar{x}\forall\bar{y}\delta(\bar{x},\bar{y})$, where $\delta$ is a positive first-order formula using fixed distances as predicates (called a distance formula in [20]).
-On every graph $G$ a formula $\delta(\bar{x},\bar{y})$ defines a bipartite graph $G_{\delta}$ on the vertex $V(G)^{|\bar{x}|}\times V(G)^{|\bar{y}|}$, where two elements $\bar{u}$ and $\bar{v}$ are connected by an edge if the formula $\delta(\bar{u},\bar{v})$ is true in $G$.
+On every graph $G$, a formula $\delta(\bar{x},\bar{y})$ defines a bipartite graph $G_{\delta}$ on the vertex set $V(G)^{|\bar{x}|}\times V(G)^{|\bar{y}|}$, where two elements $\bar{u}$ and $\bar{v}$ are connected by an edge if the formula $\delta(\bar{u},\bar{v})$ is true in $G$.
 The semi-ladder index of $\delta$ on $G$ is defined as the semi-ladder index of $G_{\delta}$.
@@
 Assume that $G_{\delta}$ has bounded semi-ladder index $\ell$ for some fixed $\ell\in\mathbb{N}$ on all graphs $G\in\mathscr{C}$.
-Then there is an algorithm that solves the domination-type problem $\delta$ on graphs $G$ from $\mathscr{C}$ in time $f(\ell,|\varphi|)\cdot\|G\|$.
-Annotated Partial Domination can be written with the following first order formula:
-$\varphi(k,d)=\exists x_{1}\ldots\exists x_{k}$ $\exists y_{1}\ldots\exists y_{d}\forall z\leavevmode\nobreak\ \leavevmode\nobreak\ \Big{(}\bigwedge_{1\leq i\leq k}\neg F(x_{i})\Big{)}\wedge\Big{(}\bigwedge_{1\leq i\leq d}B(y_{i})\Big{)}\wedge$
+Then there is an algorithm that solves the domination-type problem $\delta$ on graphs $G$ from $\mathscr{C}$ in time $f(\ell,|\delta|)\cdot\|G\|$.
+Annotated Partial Domination can be written with the following first-order formula:
+$\varphi_{k,d}=\exists x_{1}\ldots\exists x_{k}$ $\exists y_{1}\ldots\exists y_{d}\forall z\leavevmode\nobreak\ \leavevmode\nobreak\ \Big{(}\bigwedge_{1\leq i\leq k}\neg F(x_{i})\Big{)}\wedge\Big{(}\bigwedge_{1\leq i\leq d}B(y_{i})\Big{)}\wedge$
 $\Big{(}R(z)\rightarrow\Big{(}\bigvee_{1\leq i\leq k}\mathrm{dist}(z,x_{i})=0\vee\bigvee_{1\leq i\leq d}\mathrm{dist}(z,y_{i})\leq 1\Big{)}\Big{)}$
-However, $\varphi$ is not a distance formula as it uses negations. This can be circumvented by inverting the predicate $F^{\prime}=V(G)-F$ and $R^{\prime}=V(G)-R$. We then obtain the following formula
+However, $\varphi$ is not a distance formula as it uses negations.
+This can be circumvented by inverting the predicates, i.e.
+$F^{\prime}=V(G)-F$ and $R^{\prime}=V(G)-R$.
+We then obtain the following formula
 $\delta_{k,d}=\exists x_{1}\ldots\exists x_{k}$ $\exists y_{1}\ldots\exists y_{d}\forall z\leavevmode\nobreak\ \leavevmode\nobreak\ \Big{(}\bigwedge_{1\leq i\leq k}F^{\prime}(x_{i})\Big{)}\wedge\Big{(}\bigwedge_{1\leq i\leq d}B(y_{i})\Big{)}\wedge$
@@
 Then Annotated Partial Domination can be solved in time $f(d,k,\ell)\cdot\|G\|$ for all $G\in\mathscr{C}$ for a computable function $f$.
-While the semi-ladder index is currently one of the most general graph parameter enabling FPT evaluation of $\delta_{k,d}$, some other parameter could have this property.
+While the semi-ladder index is currently one of the most general graph parameters enabling FPT evaluation of $\delta_{k,d}$, some other parameter could have this property.
 Property 1.
-A graph parameter $t(\cdot)$ has 1 if for every graph $G$ Annotated Partial Domination can be solved in time $f(d,k,t(G))\cdot|G|^{\mathcal{O}(1)}$ for a computable function $f$.
-As previously discussed, examples of parameters with 1 include treewidth, maximum degree, Hadwiger number (i.e.
+A graph parameter $t(\cdot)$ has ˜1 if for every graph $G$ Annotated Partial Domination can be solved in time $f(d,k,t(G))\cdot|G|^{\mathcal{O}(1)}$ for a computable function $f$.
+As previously discussed, examples of parameters with ˜1 include treewidth, maximum degree, Hadwiger number (i.e.
 largest clique minor), degeneracy, largest excluded bi-clique, as they all imply bounded semi-ladder index.
@@
 As mentioned, we want to solve Annotated Partial Domination on slightly modified graphs.
-See Definition B.6 for the precise definition of bag graphs.
-We therefore want a parameter that, if bounded on the graph, bounds a parameter with 1 on bag graphs.
+See Definition˜B.6 for the precise definition of bag graphs.
+We therefore want a parameter that, if bounded on the graph, bounds a parameter with ˜1 on bag graphs.
 Property 2.
-A graph parameter $t_{2}(\cdot)$ has 2 if there is a parameter $t_{1}(\cdot)$ satisfying 1 such that for every graph $G$, and integers $q,k,d$, every $(q,k)$-unbreakable tree decomposition $(T,\mathsf{bag})$, vertex $x\in T$, and set $S\subseteq\mathsf{cone}(x)$ we have that $t_{1}(\mathsf{bgraph}_{S}(x))\leq f(q,k,d,t_{2}(G))$.
-In the rest of the paper, we precisely reference parameters having 1 or 2. In particular, the semi-ladder index also satisfies 2, see Lemma B.11.
+A graph parameter $t_{2}(\cdot)$ has ˜2 if there is a parameter $t_{1}(\cdot)$ satisfying ˜1 such that for every graph $G$, and integers $q,k,d$, every $(q,k)$-unbreakable tree decomposition $(T,\mathsf{bag})$, vertex $x\in T$, and set $S\subseteq\mathsf{cone}(x)$ we have that $t_{1}(\mathsf{bgraph}_{S}(x))\leq f(q,k,d,t_{2}(G))$.
+In the rest of the paper, we precisely reference parameters having ˜1 or ˜2. In particular, the semi-ladder index also satisfies ˜2, see Lemma˜B.11.
 4 Algorithm on unbreakable graphs based on Annotated Partial Domination
 As sketched above, we first solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graphs.
-Our algorithm runs efficiently on all graph classes for which we can efficiently solve the Annotated Partial Domination problem, that is, for classes on which a parameter $t(\cdot)$ satisfying 1 is bounded.
+Our algorithm runs efficiently on all graph classes for which we can efficiently solve the Annotated Partial Domination problem, that is, for classes on which a parameter $t(\cdot)$ satisfying ˜1 is bounded.
 In the following, $q,k$, and $d$ will always be non-negative integers, and we will not always quantify them for readability.
-We will also tacitly assume that our graphs have at least $12q+$ 1 vertices.
-Then, whenever we delete a set $S$ of at most $k$ vertices there will be a unique connected component $C_{0}$ of $G-S$ with more than $q$ vertices that we call the large connected component.
+We will also tacitly assume that our graphs have at least $2q+1$ vertices.
+Then, whenever we delete a set $S$ of at most $k$ vertices there will be a unique connected component $C_{0}$ of GSG-S with more than $q$ vertices that we call the large connected component.
 When a graph has less vertices, we can solve all our problems by brute-force.
 Lemma 4.1.
-Let $G$ be a $(q,k)$-unbreakable graph, and let $v\in V(G)$. Then $G-v$ is $(q,k-1)$-unbreakable.
+Let $G$ be a $(q,k)$-unbreakable graph, and let $v\in V(G)$. Then GvG-v is (q,k1)(q,k-1)-unbreakable.
 Proof 4.2.
-Assume that there is a connected component of $G-v$ that is not $(q,k-1)$-unbreakable.
-Then, there is a separation $(A,B)$ of $G-v$ such that $|A|\geq q$, $|B|\geq q$ and $|A\cap B|\leq k-1$.
-By adding the vertex $v$ to the separator, we obtain a separation $(A^{\prime},B^{\prime})$ of $G$ such that $|A^{\prime}|>q$, $|B^{\prime}|>q$ and $|A^{\prime}\cap B^{\prime}|\leq k$ contracting that $G$ is $(q,k)$-unbreakable.
+Assume that GvG-v that is not (q,k1)(q,k-1)-unbreakable.
+Then, there is a separation $(A,B)$ of GvG-v such that $|A|\geq q$, $|B|\geq q$ and $|A\cap B|\leq k-1$.
+By adding the vertex $v$ to the separator, we obtain a separation $(A^{\prime},B^{\prime})$ of $G$ such that $|A^{\prime}|>q$, $|B^{\prime}|>q$ and $|A^{\prime}\cap B^{\prime}|\leq k$ contradicting that $G$ is $(q,k)$-unbreakable.
 Lemma 4.3.
@@
 Let $A=S^{\prime}\cup C_{0}$, where $C_{0}$ is the large connected component of $G-S^{\prime}$ and $B=S^{\prime}\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of $G-S^{\prime}$.
-Then $(A,B)$ is a separation of order at most $k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable.
+Then, $(A,B)$ is a separation of order at most $k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable.
 As $G$ is a positive instance of Dominated Cluster Deletion, $C_{0}$ can be dominated by a set $D_{0}$ of size at most $d$.
@@
   • $S$ contains exactly the vertices of $S^{\prime}$ that have at least one neighbor in $C^{\prime}_{0}$ and at least one neighbor in $G-N[C^{\prime}_{0}]$, where $C^{\prime}_{0}$ is the large connected component of $G-S^{\prime}$.
-Note that $S$ may be the empty set in the third item above if $G-S^{\prime}$ has only one connected component. Figure 3 shows examples of skeletons for dominated cluster deletion.
+Note that $S$ may be the empty set in the third item above if $G-S^{\prime}$ has only one connected component. Figure˜3 shows examples of skeletons for dominated cluster deletion.
 Figure 3: Different skeletons for the dominated cluster deletion problem for parameter $k=4$ and $d=1$ on a fixed (14,4)-unbreakable graph.
@@
 Proof 4.7.
-Assume that the Case 1 is not satisfied.
-Meaning that no vertices from the skeleton are in the dominating set $X$.
+Assume that Case 1 is not satisfied so that no vertices from the skeleton are in the dominating set $X$.
 Denote by $C^{\prime}_{0}$ the large connected component of $G-S^{\prime}$ for a set $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, hence, $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$.
-Let $C_{0}$ be the large connected component of $G-S$.
+Let $C_{0}$ be the large connected component of GSG-S.
 Note that it contains $C^{\prime}_{0}$.
-Let $A=S\cup C_{0}$ and $B=S\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of $G-S$.
+Let $A=S\cup C_{0}$ and $B=S\cup\bigcup_{1\leq j\leq m}C_{j}$, where $C_{1},\ldots,C_{m}$ are the other connected components of GSG-S.
 Then $(A,B)$ is a separation of order at most $|S|\leq k$, hence $|B|\leq q$, as $G$ is $(q,k)$-unbreakable.
 Pick an arbitrary element $v\in S$.
-By definition of $S$, $v$ has a neighbor in $C_{0}$ and in a second connected component, say $C_{1}$ of $G-S$.
-This neighbor must be a vertex $w$ of degree at most $q$ as its neighborhood lies in $G-C_{0}\subseteq B$.
+By definition of $S$, $v$ has a neighbor in $C_{0}$ and a neighbor $w$ in a second connected component, say $C_{1}$ of GSG-S.
+Hence, $w$ must be a vertex of degree at most $q$ as its neighborhood lies in $G-C_{0}\subseteq B$.
 If $w$ is in $X$, then this vertex $v$ satisfies Case 2.
@@
 Hence, similarly to $w$, we have that $x$ has degree at most $q$, and we are in Case 3.
-Lemma 4.6 gives rise to a bounded search tree branching algorithm that branches over the possible choices for a vertex from $S$ given by the conditions of the lemma.
+Lemma˜4.6 gives rise to a bounded search tree branching algorithm that branches over the possible choices for a vertex from $S$ given by the conditions of the lemma.
 Definition 4.8.
@@
 Second, it computes the set $X\cup Y\cup Z$, where $Y$ is the set of all vertices in the neighborhood $N(x)$ of a vertex $x\in X$ with $\deg(x)\leq q$, and the set $Z$ is the set of all vertices in the neighborhood $N(y)$ of a vertex $y\in Y$ where $\deg(y)\leq q$.
-Third, call the algorithm with parameters $q,k-1,d$ on $G-v$ for every $v$ in $X\cup Y\cup Z$ (and get a family $\mathcal{F}_{v}$).
+Third, call the algorithm with parameters q,k1,dq,k-1,d on GvG-v for every $v$ in $X\cup Y\cup Z$ (and get a family $\mathcal{F}_{v}$).
 Finally, for every $v$ in $X\cup Y\cup Z$ add $v$ to every set in $\mathcal{F}_{v}$, and return the empty set and the union of all the families $\mathcal{F}_{v}$.
-Note that in the above definition, if a family $\mathcal{F}_{v}$ is empty, adding $v$ to every set still results in an empty family (e.g.
-in the case where $G-v$ does not have a $q+d$ dominating set).
-To not be confused with $\mathcal{F}_{v}$ having the empty set as one of the family member (e.g.
-for branches of the algorithm that stops before reaching $k=0$).
 Algorithm 1 Skeleton-algorithm
@@
 13:end procedure
-We now show that the skeleton algorithm 1) runs efficiently and outputs a small family of sets.
-This is proved in Lemma 4.9, and 2) if the given graph was a positive instance to Dominated Cluster Deletion, then the skeleton algorithm outputs all possible skeletons; see Lemma 4.11.
+Note that in the above definition, if a family $\mathcal{F}_{v}$ is empty, adding $v$ to every set still results in an empty family (e.g.
+in the case where GvG-v does not have a $q+d$ dominating set).
+Do not be confused with $\mathcal{F}_{v}$ having the empty set as one of the family members (e.g.
+for branches of the algorithm that stops before reaching $k=0$).
+We now show that the skeleton algorithm 1) runs efficiently and outputs a small family of sets which is proved in Lemma˜4.9, and 2) if the given graph was a positive instance to Dominated Cluster Deletion, then the skeleton algorithm outputs all possible skeletons; see Lemma˜4.11.
 Lemma 4.9.
-Let $t(\cdot)$ be a graph parameter satisfying 1, then the skeleton-algorithm runs in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, and outputs a family of at most $g(q,k,d)$ sets where $f$ and $g$ are computable functions.
+Let $t(\cdot)$ be a graph parameter satisfying ˜1, then the skeleton-algorithm runs in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, and outputs a family of at most $g(q,k,d)$ sets where $f$ and $g$ are computable functions.
 Proof 4.10.
-Computing $X,Y$ and $Z$ takes time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$ thanks to the fact that $t(\cdot)$ satisfies 1.
-We then branch over $(q+d)q^{2}$ many choices for $v$, and recurse at most $k$ times.
+Computing $X,Y$ and $Z$ takes time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$ thanks to the fact that $t(\cdot)$ satisfies ˜1.
+We then branch over $(q+d)q^{2}$ many choices for $v$, and the recursion has depth at most $k$.
 Lemma 4.11.
@@
 If $k=0$, we simply output the empty set.
-Otherwise, by Lemma 4.3 there is such a set $X$, and by Lemma 4.6 at least one element $v$ of $X$ is in $S$.
-For the correct choices of $v$, $G-v$ is a positive instance for Dominated Cluster Deletion with parameter $k-1,d$, and by Lemma 4.1, $G-v$ is $(q,k-1)$-unbreakable, so the recursive call to Skeleton produces the expected sets by induction.
+Otherwise, by Lemma˜4.3 there is such a set $X$, and by Lemma˜4.6 at least one element $v$ of $X$ is in $S$.
+For the correct choices of $v$, GvG-v is a positive instance for Dominated Cluster Deletion with parameter k1,dk-1,d, and by Lemma˜4.1, GvG-v is (q,k1)(q,k-1)-unbreakable, so the recursive call to Skeleton produces the expected sets by induction.
 It remains to verify that $S$ can indeed be enlarged to a set $S^{\prime}$ of size at most $k$ with the properties.
 Lemma 4.13.
-Given a $(q,k)$-unbreakable graph $G$ and a set $S$ of at most $k$ vertices, we can test in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, whether $S$ is a skeleton for Dominated Cluster Deletion with parameters $k$ and $d$, and $t(\cdot)$ satisfies 1.
+Given a $(q,k)$-unbreakable graph $G$ and a set $S$ of at most $k$ vertices, we can test in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, whether $S$ is a skeleton for Dominated Cluster Deletion with parameters $k$ and $d$ assuming $t(\cdot)$ satisfies ˜1.
 Proof 4.14.
-We need to check the existence of a set $S^{\prime}\supseteq S$ such that:
+We need to check the existence of a set $S^{\prime}\supseteq S$ with $|S^{\prime}|\leq k$ such that:
   • every connected component of $G-S^{\prime}$ has domination number at most $d$, and
   • $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$ where $C^{\prime}_{0}$ is the large connected component of $G-S^{\prime}$.
-Note that by these conditions the large connected component $C_{0}$ of $G-S$ is separated from the small connected components by $S$. What remains to be done is hence the following.
+Note that by these conditions the large connected component $C_{0}$ of GSG-S is separated from the small connected components by $S$. Let $s=|S|$, what remains to be done is hence the following.
   1. 1.
     Find an optimal solution for $G-S-C_{0}$.
-This can be done in time $f(k,d)$, because this subgraph has size at most $q$ and yields a value $k^{\prime\prime}$ of vertices that need to be deleted.
-If $k^{\prime\prime}>k-k^{\prime}$ we reject the instance (in this branch).
+This can be done in time $f(k,d)$, because this subgraph has size at most $q$ and yields a value $k^{\prime}$ of vertices that need to be deleted.
+If $k^{\prime}>k-s$, we conclude that this $S$ is not a skeleton because there is no possible $S^{\prime}$.
   2. 2.
-    Test whether an additional set $W$ of $k^{\prime\prime\prime}=k-k^{\prime}-k^{\prime\prime}$ vertices can be deleted from $C_{0}$ such that $G[C_{0}-W]$ can be dominated by $d$ vertices.
+    Test whether an additional set $W$ of size $k^{\prime\prime}=k-s-k^{\prime}$ can be deleted from $C_{0}$ such that $G[C_{0}-W]$ can be dominated by $d$ vertices.
 Note that by assumption on $S$ and $S^{\prime}$ the deletion of $W$ will not cause $C_{0}$ to break into multiple further connected components.
-Hence, the remaining problem is an instance of Partial Domination with parameter $k^{\prime\prime\prime}$.
+Hence, we simply check whether $C_{0}$ is a positive instance of Annotated Partial Domination with parameter $k^{\prime\prime}$, where $R=B=C_{0}$, and $F=\emptyset$.
 If $S$ is a skeleton, such a set exists (as $S^{\prime}\cap C_{0}$ is a candidate) and hence, if $G$ is a positive instance of Dominated Cluster Deletion a positive solution will be found.
-Since $t(\cdot)$ is a graph parameter satisfying 1, we can test the existence of such $W$ in time $f(k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$.
-If no solution is found we reject the instance (in this branch).
-The combination of Lemmas 4.6, 4.9, 4.11 and 4.13 yields the following theorem.
+Since $t(\cdot)$ is a graph parameter satisfying ˜1, we can test the existence of such $W$ in time $f(k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$.
+If no solution is found, then this $S$ is not a skeleton.
+The combination of Lemmas˜4.6, 4.9, 4.11 and 4.13 yields the following theorem.
 Theorem 4.15.
-Dominated Cluster Deletion on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
+Dominated Cluster Deletion on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.
 4.2 Elimination Distance to Dominated Clusters
@@
 Of course, the notion of skeletons has to be adapted, however, they behave very similarly on unbreakable graphs.
-First let’s properly define Elimination Distance to Dominated Clusters.
+First, let us properly define Elimination Distance to Dominated Clusters.
 Definition 4.16.
@@
   • $G$ is not connected and every connected component of $G$ has elimination distance $k$ to $\mathscr{C}$, or
-  • $G$ is connected and there is a vertex $v$ such that $G-v$ has elimination distance $k-1$ to $\mathscr{C}$, or
+  • $G$ is connected and there is a vertex $v$ such that GvG-v has elimination distance k1k-1 to $\mathscr{C}$, or
   • $G\in\mathscr{C}$ and $k=0$.
-In our case we measure, for every $d$, the elimination distance to the class $\mathscr{C}_{d}$ of $d$-dominated clusters i.e., graphs where every connected component has a $d$-dominating set.
+In our case we measure, for every $d$, the elimination distance to the class $\mathscr{C}_{d}$ of $d$-dominated clusters, i.e.
+graphs where every connected component has a $d$-dominating set.
 We interchangeably write “elimination distance $k$ to $d$-dominated cluster”, and “a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters”.
@@
 While this set is of unbounded size, it has a nice tree structure that is much more useful for our algorithms.
-This characterization is pretty standard, see the definition of elimination order [9, Proposition 4.3]
+This characterization is pretty standard, see the definition of elimination order [10, Proposition 4.3]
 For a rooted tree $T$, we write $\preccurlyeq_{T}$ for the ancestor relation, that is, $x\preccurlyeq_{T}y$ if $x$ lies on the unique path from $y$ to the root of $T$.
@@
 Let $G$ be a graph and $S\subseteq V(G)$.
-We say that $S$ is tree-structured if there is a rooted tree $T$ and a mapping $\lambda:S\rightarrow V(T)$ such that every path in $G$ between two vertices $u,v$ of $S$ must contain a common ancestor in $T$ of $\lambda(u)$ and $\lambda(v)$.
+We say that $S$ is tree-structured if there is a rooted tree $T$ and a bijection $\lambda:S\rightarrow V(T)$ such that every path in $G$ between two vertices $u,v$ of $S$ must contain a common ancestor in $T$ of $\lambda(u)$ and $\lambda(v)$.
 The tree $T$ is called an elimination tree for $S$.
-The elimination depth of a tree-structured set with respect to the tree $T$ is the depth of $T$.
-The elimination depth of $S$ is the minimum elimination depth of $S$ over all elimination trees for $S$.
+The elimination depth of a tree-structured set with respect to the tree $T$ is the depth of $T$.The elimination depth of $S$ is the minimum elimination depth of $S$ over all elimination trees for $S$.
 Remark 4.18.
-A graph $G$ is a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters if and only if there is a tree-structured set $S$ with elimination depth $k$ such that every connected component of $G-S$ has a $d$-dominating set.
+A graph $G$ is a positive instance for $(k,d)$-Elimination Distance to Dominated Clusters if and only if there is a tree-structured set $S$ with elimination depth $k$ such that every connected component of GSG-S has a $d$-dominating set.
 Lemma 4.19.
-Let $G$ be a $(q,k)$-unbreakable graph. Let $S$ be a tree-structured set with an elimination tree $T$ of depth at most $k$. Then
+Let $G$ be a $(q,k)$-unbreakable graph with at least $3q(k+q)$ vertices. Let $S$ be a tree-structured set with an elimination tree $T$ of depth at most $k$. Then
   1. 1.
-    the largest connected component $C_{0}$ of $G-S$ is uniquely determined, and
+    the largest connected component $C_{0}$ of GSG-S is uniquely determined, and
   2. 2.
-    for every component $C$ of $G-S$ there is a subset $S(C)\subseteq S$ of size at most $k$ that separates $C$ from the rest, that is, $G-S(C)$ has $C$ as one of its components.
+    for every component $C$ of GSG-S there is a subset $S(C)\subseteq S$ of size at most $k$ that separates $C$ from the rest, that is, GS(C)G-S(C) has $C$ as one of its components.
 Proof 4.20.
-Let $C$ be a connected component of $G-S$.
+Let $C$ be a connected component of GSG-S.
 As $S$ is tree-structured by $T$, $C$ is not adjacent to two $\preccurlyeq_{T}$-incomparable elements, that is, the set $S(C)$ of neighbors of $C$ in $S$ consists of $\preccurlyeq_{T}$-comparable elements.
-$C$ is a connected component both of $G-S$ and of $G-S(C)$, as all neighbors of $C$ in $S$ are collected in $S(C)$.
-Assume now that there are two connected components $C_{0},C_{1}$ of $G-S$ of size greater than $q$.
-However, $C_{0}$ and $C_{1}$ are separated by $S(C_{0})$ of size at most $k$, which contradicts the $(q,k)$-unbreakability of $G$ (assuming that $G$ has at least $12q+$ 1 vertices).
-Hence, the largest connected component $C_{0}$ of $G-S$ is the unique connected component of size greater than $q$.
-Lemma 4.21 (Analog of Lemma 4.3).
+$C$ is a connected component both of GSG-S and of GS(C)G-S(C), as all neighbors of $C$ in $S$ are collected in $S(C)$.
+Assume now that there are two connected components $C_{0},C_{1}$ of GSG-S of size greater than $q$.
+However, $C_{0}$ and $C_{1}$ are separated by $S(C_{0})$ of size at most $k$, which contradicts the $(q,k)$-unbreakability of $G$.
+So there can only be one connected component of size greater than $q$.
+We now show that there must be at least one such component.
+First note that there cannot be more than $3q$ connected component in GSG-S.
+Otherwise, there would be a balance separator of $T$ of size $1$, by deleting this vertex and its ancestors, we delete at most $k$ vertices and obtain a separation with both sides containing at least $q$ components of GSG-S, contradicting the unbreakability of $G$.
+Last, as there are only $3q$ components $C$ and $|S(C)|\leq k$ we have that $|S|\leq 3qk$ hence $|G-S|\geq 3q^{2}$ so on of the component must contain at least $q$ vertices.
+Lemma 4.21 (Analog of Lemma˜4.3).
 Let $G$ be a $(q,k)$-unbreakable graph. If $G$ is a positive instance of Elimination Distance to Dominated Clusters with parameters $k$ and $d$, then $G$ has a dominating set of size at most $q+d$.
-The proof can be taken verbatim from Lemma 4.3 with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
+The proof can be taken verbatim from Lemma˜4.3 with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
 Definition 4.22.
@@
   • $S$ contains exactly the vertices of $S^{\prime}$ that have a neighbor in $C^{\prime}_{0}$ and a second neighbor in $G-N[C^{\prime}_{0}]$ where $C^{\prime}_{0}$ is the large connected component of $G-S^{\prime}$ (the next lemma shows that this connected component is uniquely determined even though $S^{\prime}$ can be larger than $k$).
-The following observation follows immediately from Lemma 4.19. {observation} Every skeleton has size at most $k$.
+The following observation follows immediately from Lemma˜4.19. {observation} Every skeleton has size at most $k$.
 Figure 4: Example of a skeleton for elimination distance to dominated clusters for parameter $k=3$ and $d=1$ on a (12,3)-unbreakable graph (left).
 The skeleton $S$ is in red, the witness $S^{\prime}$ also contains the crossed-out vertices.
-A tree structure for $S^{\prime}$ of elimination depth $3$ is on the right.
-Now the analog of Lemma 4.6 also holds for skeletons for elimination distance to dominated clusters. The proof can be taken verbatim with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
+A tree structure for $S^{\prime}$ of elimination depth 33 is on the right.
+Now the analog of Lemma˜4.6 also holds for skeletons for elimination distance to dominated clusters. The proof can be taken verbatim with $S^{\prime}$ replaced by $S^{\prime}(C_{0})$.
 Lemma 4.23.
@@
 Hence, the algorithm to guess a skeleton $S$ can be carried out in exactly the same way as for Dominated Cluster Deletion.
-We then guess a linear order on $S$ (which takes time at most $\mathcal{O}(k!)$).
-In the final step when we aim to extend a fixed skeleton $S$ to a tree-structured solution $S^{\prime}$, the only difference is that we have to solve the base problem Elimination Distance to Dominated Clusters for the small part $G-S-C^{\prime}_{0}$, which again can be simply solved by brute-force.
-Note that because $G-S-C^{\prime}_{0}$ is small we can also find an elimination tree $T$ for $S^{\prime}$ and embed $S$ according to the originally guessed order.
+We then guess a tree structure on $S$ (which takes time at most $\mathcal{O}(k!)$).
+Note that these vertices being adjacent to the large connected component, the tree structure is a linear order.
+In the final step, when we aim to extend a fixed skeleton $S$ to a tree-structured solution $S^{\prime}$, the only difference is that we have to solve the base problem Elimination Distance to Dominated Clusters for the small part $G-S-C^{\prime}_{0}$, which again can be simply solved by brute-force.
+Note that because $G-S-C^{\prime}_{0}$ is small, we can also find an elimination tree $T$ for $S^{\prime}$ and embed $S$ according to the originally guessed order.
 We conclude the main theorem of this section.
-The full algorithm is presented in the appendix.
+The full algorithm for the annotated version is presented in the appendix.
 Theorem 4.24.
-Elimination Distance to Dominated Clusters on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
+Elimination Distance to Dominated Clusters on $(q,k)$-unbreakable graphs can be solved in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.
 4.3 Skeletons for general graphs
-In this section, we show that by parameterizing only by $q,k$, and $d$ we can compute the skeleton of unbreakable graphs in time $f(q,k,d)\cdot|G|^{\mathcal{O}(d)}$.
+In this section, we show that by parameterizing only by $q,k$, and $d$, we can compute the skeleton of unbreakable graphs in time $f(q,k,d)\cdot|G|^{\mathcal{O}(d)}$.
 This construction follows the work of Bentert et al.
@@
 There is an algorithm that, given a $(q,k)$-unbreakable graph $G$, computes all skeletons $S$ for dominated cluster deletion and for elimination distance to dominated clusters for parameters $k$ and $d$ in time $f(q,k,d)\cdot n^{\mathcal{O}(d)}$.
-This lemma resembles and can be proved analogously to [3, Lemma 7].
+This lemma resembles and can be proved analogously to [4, Lemma 7].
 Proof 4.26.
-Assume the existence of a solution $S^{\prime}$ and let $S$ be it skeleton.
+Assume the existence of a solution $S^{\prime}$ and let $S$ be its skeleton.
 We guess a set $D$ of at most $d$ vertices in time $O(n^{d})$.
@@
 We then define $S_{1}=G-N[D]$ and $S_{2}=\{v\in N(u)\leavevmode\nobreak\ |\leavevmode\nobreak\ u\in S_{1}\wedge\deg(u)\leq q\}$.
-If either $|S_{1}|$ or $|S_{2}|$ is bigger than $q$, we terminate this branch.
+If either $|S_{1}|>q$ or $|S_{2}|>q+kq$, we conclude that this $D$ is not a good candidate for the dominating set of the large connected component and terminate this branch.
 Otherwise, we now claim that $S\subseteq S_{2}$.
@@
 Observe that $C^{\prime}_{0}\subseteq N[D]\subseteq N[C^{\prime}_{0}]$.
-Also, a vertex of $G-N[C^{\prime}_{0}]$ has no neighbor in $C^{\prime}_{0}$, and therefore only has degree at mots $q$.
-Finally, remember that any vertex $v$ of $S$ must have neighbor $u$ that is not in $N[C^{\prime}_{0}]$ and hence not in $N[G]$.
+A vertex of $G-N[C^{\prime}_{0}]$ has no neighbor in $C^{\prime}_{0}$, and therefore only has degree at most $q$.
+Finally, remember that any vertex $v$ of $S$ must have a neighbor $u$ that is not in $N[C^{\prime}_{0}]$ and hence not in $N[D]$.
 So $u$ is in $S_{1}$ and $u$ has degree at most $q$.
 So $v$ belongs to $S_{2}$.
+Regarding the cardinality constraints.
+Note that at most $q$ vertices are in $G-C^{\prime}_{0}$ and hence at most $q$ are in $G-N[D]=S_{1}$.
+Also, the only vertices of $S_{2}$ that are in $C_{0}$ must be neighbors of a vertex of $S^{\prime}$ of degree at most $q$, hence $|S_{2}\cap C_{0}|\leq kq$, and since $|G-C_{0}|\leq q$, we get $|S_{2}|\leq q+kq$.
 We conclude that $S\subseteq S_{2}$. So we can guess $S$ in time $f(q,k,d)\cdot n^{\mathcal{O}(d)}$.
-Note that Lemma 4.25 applies to skeletons for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters.
+Note that Lemma˜4.25 applies to skeletons for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters.
 Furthermore, the dominating set $D$ for the large connected component is computed (by brute-force).
@@
 Definition 5.1 (Annotated Dominated Cluster Deletion).
-Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Dominated Cluster Deletion problem is to find at most $k$ vertices $S\subseteq G-F$ such that every connected component $C$ of $G-S$ has a Red-Blue dominating set of size at most $d$ (i.e.
+Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Dominated Cluster Deletion problem is to find at most $k$ vertices $S\subseteq G-F$ such that every connected component $C$ of GSG-S has a Red-Blue dominating set of size at most $d$ (i.e.
 $d$ Blue vertices dominating all Red vertices).
-The next proposition states that this problem can be solved on specific instances which we call bag graphs.
+The next proposition states that this problem can be solved on instances we call bag graphs.
 Definition 5.2.
-A graph $G$ is called a $(q,k,d)$-bag graph, if the vertices of $G$ can be partition into two vertex sets $I,E$ called respectively interior vertices and exterior vertices, such that:
-  • a connected component of $G[E]$ has at most $12d+$ 1 vertices,
-  • a connected component of $G[E]$ has at most $q$ neighbors in $I$, and
-  • for every separation $L,R$ of $G$ of order $k$ such that $L\cap R\subseteq I$, we have that either $|L\cap I|\leq q$, or $|R\cap I|\leq q$.
+A graph $G$ is called a $(q,k,d)$-bag graph, if the vertices of $G$ can be partitioned into two vertex sets $I,E$ called respectively interior vertices and exterior vertices, such that:
+  • every connected component of $G[E]$ has at most $2d+1$ vertices,
+  • every connected component of $G[E]$ has at most $q$ neighbors in $I$, and
+  • for every separation $L,R$ of $G$ of order $k$ such that $L\cap R\subseteq I$, we have that $|L\cap I|\leq q$ or $|R\cap I|\leq q$.
 Proposition 5.3.
-There is an algorithm such that, given a $(q,k,d)$-bag graph $G=(I,E)$, and vertex sets $F,R,B$ of $G$ satisfying $E\subseteq F$ and $|E\cap B|\leq qd$, the algorithm decides whether $G$ is a positive instance to $(k,d)$-Annotated Dominated Cluster Deletion, in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
-We prove Proposition 5.3 in Appendix A. Using this, we then prove the following statement via dynamic programming on an unbreakable tree decomposition in Appendix B.
+There is an algorithm such that, given a $(q,k,d)$-bag graph $G=(I,E)$, and vertex sets $F,R,B$ of $G$ satisfying $E\subseteq F$ and $|E\cap B|\leq qd$, the algorithm decides whether $G$ is a positive instance to $(k,d)$-Annotated Dominated Cluster Deletion, in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.
+We prove Proposition˜5.3 in Appendix˜A. Using this, we then prove the following statement via dynamic programming on an unbreakable tree decomposition in Appendix˜B.
 Theorem 5.4.
-There is an algorithm solving Annotated Dominated Cluster Deletion in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying 2.
-Remember that 2 implies that, on the bag graphs that are build during the dynamic programming, some parameter with 1 is bounded, hence we can use Proposition 5.3 these bag graphs.
-Note that Theorem 5.4 implies Theorem 1.1, by setting $F=\emptyset$ and $R=B=V(G)$.
-The annotated versions for Elimination Distance to Dominated Clusters is introduced and solved in Appendix C.
+There is an algorithm solving Annotated Dominated Cluster Deletion in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying ˜2.
+Remember that ˜2 implies that some parameter with ˜1 is bounded on bag graphs built in our algorithm, so we can use Proposition˜5.3 on these bag graphs.
+Note that Theorem˜5.4 implies Theorem˜1.1, by setting $F=\emptyset$ and $R=B=V(G)$.
+The annotated version for Elimination Distance to Dominated Clusters is introduced and solved in Appendix˜C.
 6 Hardness Results
-In this section we give the details of the hardness results mentioned in the introduction.
+In this section, we give the details of the hardness results mentioned in the introduction.
 Para-NP-hardness for parameter $k+\Delta$ and for parameter $d$
-A problem is para-NP-hard with respect to a parameter if there exists an instance of the problem with a fixed parameter such that solving this instance is harder than solving any problem in NP.
+A problem is para-NP-hard with respect to a parameter if the restriction to instances whose parameter value is at most a certain constant, is itself an NP-hard problem.
 Theorem 6.1.
@@
   • $|x^{\prime}|+k^{\prime}\leq f(k)$ for a computable function $f$.
-A kernel is polynomial if $f$ is polynomial.
-[…]
-Recall that $c$ denotes the degeneracy of a graph. A graph $G$ is $c$-degenerate if every subgraph $H$ has a vertex of degree at most $c$ (in $H$).
+A kernel is polynomial if $f$ is.
+We use the AND-cross-composition technique [19] to show that Elimination Distance to Dominated Clusters does not admit a polynomial kernel unless $\textsf{NP}\subseteq\textsf{coNP/poly}$.
+If an NP-hard language $L$ AND-cross-composes into a parameterized language $Q$, then $Q$ does not admit a polynomial kernel, unless $\textsf{NP}\subseteq\textsf{coNP/poly}$.
+Since the treedepth of a disjoint union of a family of graphs is equal to the maximum over the treedepth of these graphs, the disjoint union yields an AND-cross-composition from the unparameterized version of Treedepth into the parameterized one.
+As computing the treedepth of a graph is NP-hard [32], it follows that Treedepth not admit a polynomial kernel, unless $\textsf{NP}\subseteq\textsf{coNP/poly}$.
+It remains to prove that Treedepth is also NP-complete on $2$-degenerate graphs.
+Recall that $c$ denotes the degeneracy of a graph.
+A graph $G$ is $c$-degenerate if every subgraph $H$ has a vertex of degree at most $c$ (in $H$).
 Theorem 6.7.
@@
 Our proofs combine the main tools for addressing cut problems, the decomposition theorem into unbreakable parts by Cygan et al.
-[11] with the main tools for addressing domination type problems [20].
+[12] with the main tools for addressing domination-type problems [20].
 Appendix A Algorithm on bag graphs
-This section is devoted to the proof of Proposition 5.3.
+This section is devoted to the proof of Proposition˜5.3.
 See 5.3
-It closely follows from the proof of Theorem 4.15 in Section 4.
-The main differences are that we now work on an annotated version of the problem, and that the graph might not be unbreakable, only the set of interior vertices are (see Definition 5.2).
+It closely follows from the proof of Theorem˜4.15 in Section˜4.
+The main differences are that we now work on an annotated version of the problem, and that the graph might not be unbreakable, only the set of interior vertices are (see Definition˜5.2).
 Definition A.1.
-Let $G=(I,E)$ be a $(q,k,d)$-bag graph, $F$ a set with $E\subseteq F$, $R$ and $B$ additional colors on bag vertices.
+Let $G=(I,E)$ be a $(q,k,d)$-bag graph, $F$ a set with $E\subseteq F$, $R$ and $B$ additional colors on interior vertices.
 We call a set $S$ a skeleton for annotated dominated cluster deletion with parameters $k$ and $d$ if there is a superset $S^{\prime}\supseteq S$ of vertices satisfying:
@@
   • every connected component of $G-S^{\prime}$ has a Red-Blue dominating set of size at most $d$, and
-  • $S$ contains exactly the vertices of $S^{\prime}$ that have at least one neighbor in a component $C$ of $G-S$ where $|C\cap I|\leq q$, and $C$ contains at least one Red vertex.
-We now define an operation that will be necessary to obtain a statement analogous to Lemma 4.6.
+  • $S$ contains exactly the vertices of $S^{\prime}$ that have at least one neighbor in a component $C$ of GSG-S where $|C\cap I|\leq q$, and $C$ contains at least one Red vertex.
+We now define an operation that will be necessary to obtain a statement analogous to Lemma˜4.6.
 Definition A.2.
@@
 For a set $X$, $\mathsf{sat}_{q}(X):=\bigcup\limits_{u\in X}\mathsf{sat}_{q}(u)$.
-For an integer $i$, we define the $q$-saturation at distance $i$ $\mathsf{sat}_{q}^{i}(X)$ of a vertex set $X$ as $\mathsf{sat}_{q}(X)$ if $i=1$, and $\mathsf{sat}_{q}(\mathsf{sat}_{q}^{i-1}(X))$.
-I.e.
+For an integer $i$, we define the $q$-saturation at distance $i$ $\mathsf{sat}_{q}^{i}(X)$ of a vertex set $X$ as $\mathsf{sat}_{q}(X)$ if $i=1$, and $\mathsf{sat}_{q}(\mathsf{sat}_{q}^{i-1}(X))$, i.e.
 we apply the $q$-saturation operator $i$ times.
-Intuitively from a bag vertex, we take all the bag vertices that are adjacent or connected through component vertices, but only if there are few of them (at most $q$).
+Intuitively from an interior vertex, we take all the interior vertices that are adjacent or connected through exterior vertices, but only if there are few of them (at most $q$).
 Lemma A.3.
-Let $q,k,d$ be integers, $G$ be a $(q,k)$-bag graph that is a positive instance to Annotated Dominated Cluster Deletion, then $G$ has a $3qd$ Red-Blue dominating set.
+Let $q,k,d$ be integers and $G$ be a $(q,k,d)$-bag graph that is a positive instance to Annotated Dominated Cluster Deletion, then $G$ has a $3qd$ Red-Blue dominating set.
 Proof A.4.
-Let $(I,E)$ be the partition of $G$, and $S$ be a solution i.e.
-every connected component of $G-S$ has a $d$-Red-Blue dominating set.
-Let $C_{0}$ be the only connected component of $G-S$ that contains more than $q$ vertices of $I$.
+Let $(I,E)$ be the partition of the $(q,k,d)$-bag graph $G$, and $S$ be a solution of Annotated Dominated Cluster Deletion, i.e.
+every connected component of GSG-S has a $d$-Red-Blue dominating set.
+Let $C_{0}$ be the only connected component of GSG-S that contains more than $q$ vertices of $I$.
 In $G-C_{0}$, there are at most $q$ interior vertices, which together with the $qd$ Red exterior vertices amount to at most $2qd$ Red vertices in $G-C_{0}$.
@@
 Together with the $d$ Blue vertices dominating all Red vertices of $C_{0}$, we have a $3qd$ Red-Blue dominating set.
-Lemma A.5 (analogous to Lemma 4.6).
+Lemma A.5 (analogous to Lemma˜4.6).
 Let $G=(I,E)$ be a $(q,k,d)$-bag graph with colors $F,R,B$, let $S$ be a skeleton for Annotated Dominated Cluster Deletion with parameters $k,d$, and let $X$ be a $3qd$ Red-Blue dominating set for $G$.
@@
 If there is a vertex of $S$ in $X$, we are done as $X\subseteq\mathsf{sat}_{q}^{q}(X)$.
-Otherwise, let $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, let $v$ be a vertex of $S$ and $C$ the connected component of $G-S^{\prime}$ witnessing that $v\in S$ i.e.
+Otherwise, let $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, let $v$ be a vertex of $S$ and $C$ the connected component of $G-S^{\prime}$ witnessing that $v\in S$, i.e.
 $|C\cap I|\leq q$, $C$ contains a Red vertex, and $v$ has a neighbor is $C$.
@@
 Definition A.7 (Skeleton algorithm for bag graph).
-The colored skeleton-algorithm with parameters $q,k,d$ over a bag graph $G$ returns a family of sets of vertices, each set of size at most $k$.
+The annotated skeleton-algorithm with parameters $q,k,d$ over a $(q,k,d)$-bag graph $G$ returns a family of sets of vertices, each set of size at most $k$.
 First, it computes a dominating set $X$ of size at most $3qd$ (returning the empty family if there is no such set).
 Second, it computes the set $\mathsf{sat}^{q}_{q}(X)$.
-Third, it calls the algorithm with parameters $q,k-1,d$ on $G$ (and get a family $\mathcal{F}_{0}$), and on $G-v$ for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ (and get a family $\mathcal{F}_{v}$).
+Third, it calls the algorithm with parameters q,k1,dq,k-1,d on $G$ (and get a family $\mathcal{F}_{0}$), and on GvG-v for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ (and get a family $\mathcal{F}_{v}$).
 And finally, for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ add $v$ to every set in $\mathcal{F}_{v}$, and return the union of all the families.
 Remark A.8.
-Note that for any interior vertex $v$, $G-v$ is a $(q,k-1,d)$-bag graph.
-Furthermore, if the input graph is a positive instance of Annotated Dominated Cluster Deletion, and if the algorithm guesses correctly a vertex $v$ of a solution, then $G-v$ is also a positive instance to Annotated Dominated Cluster Deletion with parameter $(k-1,d)$.
-So the algorithm can repeat until it guesses a skeleton $S$ of size at most $k^{\prime}$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
+Note that for any interior vertex $v$, GvG-v is a (q,k1,d)(q,k-1,d)-bag graph.
+Furthermore, if the input graph is a positive instance of Annotated Dominated Cluster Deletion, and if the algorithm guesses correctly a vertex $v$ of a solution, then GvG-v is also a positive instance to Annotated Dominated Cluster Deletion with parameter (k1,d)(k-1,d).
+So the algorithm can repeat until it guesses a skeleton $S$ of size at most $k^{\prime}$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.
 It remains to verify that $S$ can indeed be enlarged to a set $S^{\prime}$ of size at most $k$ with the property that every connected component of $G-S^{\prime}$ has a $d$-Red-Blue dominating set.
-Note that by definition, there is at most one component $C_{0}$ of $G-S$ that contains more than $q$ interior vertices.
+Note that by definition, there is at most one component $C_{0}$ of GSG-S that contains more than $q$ interior vertices.
 So in $G-C_{0}$, there are at most $3qd+q$ many Red vertices ($3qd$ for the exterior vertices and $q$ for the interior ones).
@@
 As we can treat indifferently two Blue vertices with the same Red neighborhood, and since there are at most $2^{O(qd)}$ possible Red neighborhoods in time $f(q,k,d)$ we can go through all possibilities to dominate the Red vertices.
-    Each such possibility yields a value $k^{\prime\prime}$ of vertices that need to be deleted.
-If $k^{\prime\prime}>k-k^{\prime}$ we reject the instance (in this branch).
-Remember that $k^{\prime}=|S|$.
+    Each such possibility yields a value $k^{\prime}$ of vertices that need to be deleted.
+If $k^{\prime}>k-s$ where $s=|S|$, we conclude that this $S$ is not a skeleton because there is no possible $S^{\prime}$.
   2. 2.
-    Test whether an additional set $W\subseteq C_{0}-F$ of $k^{\prime\prime\prime}=k-k^{\prime}-k^{\prime\prime}$ vertices that can be deleted from $C_{0}$ such that $G[C_{0}-W]$ has a $d$-Red-Blue dominating set.
-Hence, the remaining problem is an instance of Partial Red-Blue Domination with parameter $k^{\prime\prime\prime},d$.
+    Test whether an additional set $W\subseteq C_{0}-F$ of $k^{\prime\prime}=k-s-k^{\prime}$ vertices that can be deleted from $C_{0}$ such that $G[C_{0}-W]$ has a $d$-Red-Blue dominating set.
+Hence, the remaining problem is an instance of Partial Red-Blue Domination with parameter $k^{\prime\prime}$ and $d$.
 If $G$ is a positive instance of Annotated Dominated Cluster Deletion, witnessed by some set $S^{\prime}$, and if $S$ is a skeleton of $S^{\prime}$, such a set $W$ exists (as $S^{\prime}\cap C_{0}$ is a candidate).
-    If $t(\cdot)$ is a parameter satisfying 1, we can then test for the existence of such $W$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$.
-If no solution is found, we reject the instance (in this branch).
-This concludes the proof of Proposition 5.3 and therefore Appendix A.
-Appendix B Algorithm through unbreakable tree decomposition.
-In this section, we assume that we have a fix graph $G$, and $(T,\mathsf{bag})$ a $(q,k)$-unbreakable tree decomposition of $G$.
-See Section 2 for the definition of an unbreakable tree decomposition, and Theorem 2.1 ([12, Theorem 10]) for the proof of its existence.
-The goal of this section is to prove Theorem 1.1 using the fact that Annotated Dominated Cluster Deletion is FPT on bag graphs (Proposition 5.3).
+If $t(\cdot)$ is a parameter satisfying ˜1, we can then test for the existence of such $W$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$.
+If no solution is found, then this $S$ is not a skeleton.
+This concludes the proof of Proposition˜5.3 and therefore Appendix˜A.
+Appendix B Algorithm through unbreakable tree decomposition
+In this section, we assume that we have a fixed graph $G$, and a $(q,k)$-unbreakable tree decomposition $(T,\mathsf{bag})$ of $G$.
+See Section˜2 for the definition of an unbreakable tree decomposition, and Theorem˜2.1 ([13, Theorem 10]) for the proof of its existence.
+The goal of this section is to prove Theorem˜1.1 using the fact that Annotated Dominated Cluster Deletion is FPT on bag graphs (Proposition˜5.3).
 See 1.1
@@
   • $k_{A}=|S|-|S_{A}|$,
-  • for every connected component $C$ of $G-S$ either :
+  • for every connected component $C$ of GSG-S either :
     • $C\cap A=\emptyset$ and $C$ has a Red-Blue dominating set of size $d$, or
@@
 Intuitively, we mark how the set $A$ overlaps with the deleted vertices and the Blue dominating sets.
-The Blue dominating sets must dominate all Red vertices of $G-A$, and if some Red vertices of $A$ are not dominated, it is written so on $U_{A}$ (for un-dominated).
+The Blue dominating sets must dominate all Red vertices of GAG-A, and if some Red vertices of $A$ are not dominated, it is written so on $U_{A}$ (for undominated).
 Definition B.2.
@@
 Definition B.5.
-A $d$-gadget is the graph $H$ composed of $12d+$ 1 vertices $a,b_{1},\ldots,b_{d},b^{\prime}_{1}\ldots,b^{\prime}_{d}$, where for every $i\leq d$, $b_{i}$ is adjacent to $a$ and $b^{\prime}_{i}$.
+A $d$-gadget is the graph $H$ composed of $2d+1$ vertices $a,b_{1},\ldots,b_{d},b^{\prime}_{1}\ldots,b^{\prime}_{d}$, where for every $i\leq d$, $b_{i}$ is adjacent to $a$ and $b^{\prime}_{i}$.
 A $d$-gadget is plugged to a set $A$ by making every vertex of $A$ adjacent to $a$.
@@
 Let $(T,\mathsf{bag})$ be a regular $(q,k)$-unbreakable tree decomposition of a graph $G$, $S\subseteq\mathsf{cone}(x)$ a set of at most $k$ vertices, and $d$ an integer.
-We note $\mathsf{bgraph}_{S}(x)$ of a node $x\in V(T)$ any graph that is obtained from $G[\mathsf{cone}(x)-S]$ by :
+We note $\mathsf{bgraph}_{S}(x)$ of a node $x\in V(T)$ any graph that is obtained from $G[\mathsf{cone}(x)-S]$ by
   • contracting each connected component $C$ of the set $\mathsf{comp}(y)-S$ that has a neighbor in $\mathsf{adh}(y)-S$ into a $d$-gadget $H_{y}^{C}$ and plugging $H_{y}^{C}$ to the vertices of $\mathsf{bag}(x)-S$ that are neighbors of $C$, and
@@
 Lemma B.7.
-Every $\mathsf{bgraph}_{S}(x)$ is a $(q,k-|S|,d)$-bag graph.
+Every $\mathsf{bgraph}_{S}(x)$ is a (q,k|S|,d)(q,k-|S|,d)-bag graph.
 Proof B.8.
 The vertices of $\mathsf{bgraph}_{S}(x)$ are split into $I$ the vertices of $\mathsf{bag}(x)$, and $E$ the union of all the $d$-gadgets.
-Clearly, every connected component of $G-I$ is a $d$-gadget and therefore has at most $12d+$ 1 vertices.
+Clearly, every connected component of GIG-I is a $d$-gadget and therefore has at most $2d+1$ vertices.
 Furthermore, this $d$-gadget is plugged to a subset of the adhesion of $x$ or the adhesion of a child of $x$, and therefore has at most $q$ neighbors in $I$.
-For the last point, let $L,R$ be a separation of $\mathsf{bgraph}_{S}(x)$ of order at most $k-|S|$, then extend $L,R$ into $L^{\prime},R^{\prime}$ by first removing the $d$-gadget plugged to the adhesion of $x$, then adding $S$ to both $L$ and $R$, and finally replacing every $d$-gadget plugged to a subset of $\mathsf{adh}(y)$ by the component of $\mathsf{cone}(y)$ it represents.
+For the last point, let $L,R$ be a separation of $\mathsf{bgraph}_{S}(x)$ of order at most k|S|k-|S|, then extend $L,R$ into $L^{\prime},R^{\prime}$ by first removing the $d$-gadget plugged to the adhesion of $x$, then adding $S$ to both $L$ and $R$, and finally replacing every $d$-gadget plugged to a subset of $\mathsf{adh}(y)$ by the component of $\mathsf{cone}(y)$ it represents.
 We obtain that $(L^{\prime},R^{\prime})$ is a separation of order $k$ of $\mathsf{cone}(x)$, and therefore $|L^{\prime}\cap\mathsf{bag}(x)|\leq q$ or $|R^{\prime}\cap\mathsf{bag}(x)|\leq q$, hence with $I=\mathsf{bag}(x)$, either $|L\cap I|\leq q$, or $|R\cap I|\leq q$.
@@
 Furthermore, each node is the child of at most one node and each connected component of a child can be divided in at most $q$ many connected components.
-Additionally, $d$-gadgets have size at most $12d+$ 1, and $|T|\leq|G|$, hence:
+Additionally, $d$-gadgets have size at most $2d+1$, and $|T|\leq|G|$, hence:
 $\sum\limits_{x\in V(T)}|\mathsf{bgraph}(x)|\leq(2d+2)q|T|+|G|+(2d+2)q|T|\leq(4d+5)q|G|$
@@
 In the next lemma, we show that the ladder index of these bag graphs is bounded (assuming that it was already bounded in the original graph).
-This together with the fact that the semi-ladder index has 1 implies that the semi-ladder index also has 2.
+This together with the fact that the semi-ladder index has ˜1 implies that the semi-ladder index also has ˜2.
 Lemma B.11.
@@
 So $t-(q+5)+1\leq\ell$, hence $t\leq q+\ell+4$.
-We state the next proposition with arbitrary parameter $t(\cdot)$ satisfying 2. By the previous lemma, the semi-ladder index $\ell$ can replace $t(\cdot)$.
+We state the next proposition with arbitrary parameter $t(\cdot)$ satisfying ˜2. By the previous lemma, the semi-ladder index $\ell$ can replace $t(\cdot)$.
 Proposition B.13.
-Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the tree decomposition and, for each child $y$ of $x$, the $(k,d)$-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 2.
-Or alternatively in time $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$.
-Note that by Lemma B.9, the sum of the size of all bag graphs is linear in the size of the graph so that Proposition B.13 yields Theorem 5.4, and therefore Theorem 1.1.
-B.2 Algorithm for Proposition B.13
+Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the tree decomposition and, for each child $y$ of $x$, the $(k,d)$-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜2, or alternatively in time $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$.
+Note that by Lemma˜B.9, the sum of the size of all bag graphs is linear in the size of the graph so that Proposition˜B.13 yields Theorem˜5.4, and therefore Theorem˜1.1.
+B.2 Algorithm for Proposition˜B.13
 Let $G$ be a graph, $(T,\mathsf{bag})$ a $(q,k)$-unbreakable tree decomposition and $x\in V(T)$ a node of this decomposition.
@@
 Otherwise, our algorithm branches over all the realized marks of $y$.
-For each branch, considering a mark $M^{\prime}$, we check the consistency of $M^{\prime}$ with our partial solution, and if it is consistent we update our connected component labeling and our partial solution.
+For each branch, considering a mark $M^{\prime}$, we check the consistency of $M^{\prime}$ with our partial solution, and if it is consistent, we update our connected component labeling and our partial solution.
 The checks are:
@@
   • remove $N(D_{x})$ from $R$ (we mark these vertices as dominated), and
-  • for every $P\in{\cal{P}}_{x}$ plug the $(d-p)$-gadget to the vertices of $P$.
+  • for every $P\in{\cal{P}}_{x}$ plug the (dp)(d-p)-gadget to the vertices of $P$.
 We are now done building a bag graph $H$, of the form $\mathsf{bgraph}_{S}(x)$.
-We now conclude by testing whether $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion with Proposition 5.3.
+We now conclude by testing whether $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion with Proposition˜5.3.
 Claim 1 (Correctness of the algorithm).
@@
 Proof B.14.
-For the first direction, assume that $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion, then the algorithm of Proposition 5.3 produced a set $S^{\prime}$ of size $k_{x}-|S|-s$, which together with $S$ and the sets $S_{y}$ (of cumulative size $s$) selected for the children satisfies that every connected component of $\mathsf{cone}(x)-S-S^{\prime}-\bigcup\limits_{y\in\mathsf{children}(x)}S_{y}$ that does not intersect $\mathsf{adh}(x)$ has a Red-Blue dominating set of size $d$.
-For the component intersecting $\mathsf{adh}(x)$ the Red-Blue dominating set of size $d$ must contain the $d-p$ Blue vertices of the $(d-p)$-gadget; hence the component has a Red-Blue dominating set of size $p$.
+For the first direction, assume that $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion, then the algorithm of Proposition˜5.3 produced a set $S^{\prime}$ of size $k_{x}-|S|-s$, which together with $S$ and the sets $S_{y}$ (of cumulative size $s$) selected for the children satisfies that every connected component of $\mathsf{cone}(x)-S-S^{\prime}-\bigcup\limits_{y\in\mathsf{children}(x)}S_{y}$ that does not intersect $\mathsf{adh}(x)$ has a Red-Blue dominating set of size $d$.
+For the component intersecting $\mathsf{adh}(x)$ the Red-Blue dominating set of size $d$ must contain the dpd-p Blue vertices of the (dp)(d-p)-gadget; hence the component has a Red-Blue dominating set of size $p$.
 For the second direction, assume that the mark is realized, hence there is a set $\hat{S}$ on $\mathsf{cone}(x)$ witnessing it this mark is realized, together with a set of Blue vertices $D_{C}$ for every connected component of $\mathsf{cone}(x)-\hat{S}$.
@@
 Proof B.17.
-The bound on the number of leafs is straightforward.
+The bound on the number of leaves is straightforward.
 For the size, notice that whenever a White vertex has some Black children, it only increases the size of the tree to switch them (i.e.
@@
 Therefore, the maximal Black-White tree is a complete Black tree of degree $\alpha$ and of depth $\beta$, where each leaf is attached to a White path of length $n-\beta$, for which the size is at most $\alpha^{\beta}\cdot n$.
-When doing the backtracking, it is key that if a child is satisfied with the given partial solution we do not try to branch and make an impactful decision.
+When doing the backtracking, it is key that if a child is satisfied with the given partial solution, we do not try to branch and make an impactful decision.
 Otherwise, the tree of partial solutions grows too much and is not linear.
@@
 Lemma B.18.
-The branching tree of backtracking algorithm of Section B.2 can be labelled to be a Black-White tree of degree $d^{2^{\mathcal{O}(k)}}$ and of budget $(k+qd)$.
+The branching tree of the backtracking algorithm of Section˜B.2 can be labelled to be a Black-White tree of degree $d^{2^{\mathcal{O}(k)}}$ and of budget $(k+qd)$.
 Proof B.19.
-When we branch over the children we only have $d^{2^{\mathcal{O}}(k)}$ many marks to consider (see Lemma B.3). This bounds the degree of our Black-White tree.
+When we branch over the children we only have $d^{2^{\mathcal{O}}(k)}$ many marks to consider (see Lemma˜B.3). This bounds the degree of our Black-White tree.
 Furthermore, every time that we branch, this increases at least one of $|S|$, $s$, $|D|$, or $c$.
 As we terminate the algorithm as soon as $|S|+s>k$ or $|D|+c>qd$, the budget remains bounded by $k+qd$.
-Going back to the algorithm for Proposition B.13 presented in Section B.2, we just showed that the number of possible branches to consider is only a function of $q,k$, and $d$.
-And since each branch then perform a single call to Proposition 5.3, this concludes the proof of Theorem 5.4.
+Going back to the algorithm for Proposition˜B.13 presented in Section˜B.2, we just showed that the number of possible branches to consider is only a function of $q,k$, and $d$.
+And since each branch then perform a single call to Proposition˜5.3, this concludes the proof of Theorem˜5.4.
 Appendix C Annotated version for Elimination Distance to Dominated Clusters
-Similarly to Section 5, we extend Theorem 1.2 to an annotated version.
+Similarly to Section˜5, we extend Theorem˜1.2 to an annotated version.
 Definition C.1 (Annotated Elimination Distance to Dominated Clusters).
-Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Elimination Distance to Dominated Clusters problem is to find a tree-structured set $S\subseteq G-F$ of elimination depth $k$ such that every connected component $C$ of $G-S$ has a Red-Blue dominating set of size at most $d$ (i.e.
+Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Elimination Distance to Dominated Clusters problem is to find a tree-structured set $S\subseteq G-F$ of elimination depth $k$ such that every connected component $C$ of GSG-S has a Red-Blue dominating set of size at most $d$ (i.e.
 $d$ Blue vertices dominating all Red vertices).
@@
 An extended-mark is said to be realized if there is a set $S$, partitioned into sets $L^{1},\ldots,L^{k}$ of vertices such that:
-  • for every $i\leq k$ and vertices $u,v$ in $L^{i}$ every $u-v$ path must use a vertex in $L^{j<i}$,
+  • for every $i\leq k$ and vertices $u,v$ in $L^{i}$ every uvu-v path must use a vertex in $L^{j<i}$,
   • $L_{A}^{i}=L^{i}\cap A$, and
-  • for every connected component $C$ of $G-S$ either:
+  • for every connected component $C$ of GSG-S either:
     • $C\cap A=\emptyset$ and $C$ has a Red-Blue dominating set of size $d$, or
@@
 Given a graph $G$, a subset of vertices $A$, and integers $k,d$, the $(k,d)$-extended-profile of $A$ in $G$ is the list of all realized $(k,d)$-extended-marks of $A$ in $G$.
-The proof of the next observation is similar to the one of Lemma B.3.
+The proof of the next observation is similar to the one of Lemma˜B.3.
 {observation} The $(k,d)$-extended-profile of a set $A$ of size at most $q$ in a graph $G$ has size at most $(d+k)^{2^{\mathcal{O}(k)}}$.
-We are now going to work with the same definition of gadget and bag graphs (see Definitions B.5 and B.6).
+We are now going to work with the same definition of gadget and bag graphs (see Definitions˜B.5 and B.6).
 Proposition C.4.
-Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the decomposition and, for each child $y$ of $x$, the $(k,d)$-extended-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-extended-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time either $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$, or $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t(\cdot)$ is a graph parameter satisfying 2.
-Equivalently, the running time is $f(q,k,d,t_{1}(\mathsf{bgraph}(x)))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t_{1}(\cdot)$ is a graph parameter satisfying 1.
-Note that by Lemma B.9, the sum of the size of all bag graphs is linear in the size of the graph so Proposition C.4 yields our main result, Theorem 1.2.
-C.1 Algorithm for Proposition C.4
+Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the decomposition and, for each child $y$ of $x$, the $(k,d)$-extended-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-extended-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time either $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$, or $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t(\cdot)$ is a graph parameter satisfying ˜2.
+Equivalently, the running time is $f(q,k,d,t_{1}(\mathsf{bgraph}(x)))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t_{1}(\cdot)$ is a graph parameter satisfying ˜1.
+Note that by Lemma˜B.9, the sum of the size of all bag graphs is linear in the size of the graph so Proposition˜C.4 yields our main result, Theorem˜1.2.
+C.1 Algorithm for Proposition˜C.4
 Let $G$ be a graph, $(T,\mathsf{bag})$ a $(q,k)$-unbreakable tree decomposition and $x\in V(T)$ a node of this decomposition.
@@
 If the following extended-mark is realized
-${L^{i}_{y}}=\mathsf{adh}(y)\cap S^{i}\leavevmode\nobreak\ \leavevmode\nobreak\ (\text{for all }i\leq k);\leavevmode\nobreak\ \leavevmode\nobreak\ D_{y}=\mathsf{adh}(y)\cap D\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak$
+${L^{i}_{y}}=\mathsf{adh}(y)\cap S^{i}\leavevmode\nobreak\ \leavevmode\nobreak\ (\text{for all }i\leq k);\leavevmode\nobreak\ \leavevmode\nobreak\ D_{y}=\mathsf{adh}(y)\cap D\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\$
 $r^{i}=False\leavevmode\nobreak\ \leavevmode\nobreak\ (\text{for all }i\leq k);\leavevmode\nobreak\ \leavevmode\nobreak\ \bigwedge\limits_{P_{j}\in{\cal{P}}_{y}}p_{j}=0,$
@@
 Bag graphs and skeletons
-We now have a fixed set $S=\bigcup\limits_{i\leq k}S^{i}$, and similarly to Section B.2, we compute the bag graph $H_{x}=\mathsf{bgraph}_{S}(x)$.
-Similarly to Appendix A (see Lemma A.3) positive instances have a small Red-Blue dominating set.
-This transfers to the elimination distance variation, as it was the case in Lemma 4.21 for unbreakable graphs.
-So the colored version of skeleton-algorithm (see Definition A.7) can also be used here.
+We now have a fixed set $S=\bigcup\limits_{i\leq k}S^{i}$, and similarly to Section˜B.2, we compute the bag graph $H_{x}=\mathsf{bgraph}_{S}(x)$.
+Similarly to Appendix˜A (see Lemma˜A.3) positive instances have a small Red-Blue dominating set.
+This transfers to the elimination distance variation, as it was the case in Lemma˜4.21 for unbreakable graphs.
+So the colored version of skeleton-algorithm (see Definition˜A.7) can also be used here.
 Then we therefore compute every skeleton $S^{\prime}$ with parameter $k$ and $3qd$ of $H_{x}$.
@@
 We terminate this branch of the algorithm if $|S^{i}|+s_{i}>qk$, or if $S^{\prime}$ contains a vertex that lies in the adhesion of a child of $x$ (meaning that the selected mark for this child is inconsistent with the selected skeleton).
-This step can be performed in time $f(q,k,d,t_{1}(H_{x}))\cdot|H_{x}|$, where $t_{1}(\cdot)$ satisfies 1, meaning in time $f(q,k,d,t_{2}(G))\cdot|H_{x}|$, where $t_{2}(\cdot)$ satisfies 2.
+This step can be performed in time $f(q,k,d,t_{1}(H_{x}))\cdot|H_{x}|$, where $t_{1}(\cdot)$ satisfies ˜1, meaning in time $f(q,k,d,t_{2}(G))\cdot|H_{x}|$, where $t_{2}(\cdot)$ satisfies ˜2.
 Alternatively, in time $f(q,k,d)\cdot|H_{x}|^{\mathcal{O}(d)}$, in which case we also store the set $D^{\prime}$ that dominates the large connected component.
-See Lemma 4.25, which can be adapted to bag graphs.
+See Lemma˜4.25, which can be adapted to bag graphs.
 Fixing the small connected components
@@
 Partial domination in the large connected component
-We now compute a partial dominating set in the large connected component. This is achieved as for Corollary 3.4 for which the colors $F,R,B$ are defined as:
+We now compute a partial dominating set in the large connected component. This is achieved as for Corollary˜3.4 for which the colors $F,R,B$ are defined as:
   • $F$: Starts with $F$ from $G$, add every $d$-gadgets, and the adhesion of every $y$, children of $x$, together with the adhesion of $x$.
@@
 First check that in every small connected component $C_{i>0}$, every Red vertex is dominated by $D_{i}$.
-We then check that the connected components of $\mathsf{cone}(x)$ all have small Red-Blue domination sets.
-I.e.
+We then check that the connected components of $\mathsf{cone}(x)$ all have small Red-Blue domination sets, i.e.
 there could be connected components $C_{i}$, $C_{j}$ of $\mathsf{bag}(x)$ that both have a neighbor in a $d$-gadget in $H_{x}$.
@@
 Then for every $I\subset\{0,\ldots q\}$ of this partition we add up the size of $|D_{i\in I}|$, as well as the relevant $p_{j}$ from the selected marks of the children of $x$ that are adjacent to $C_{i\in I}$.
-After that, we check the consistency of the $S^{1},\ldots,S^{k}$ i.e.
+After that, we check the consistency of the $S^{1},\ldots,S^{k}$, i.e.
 no two vertices of $S^{1}$ are connected in the graphs, then no two vertices of $S^{2}$ can be connected not using $S^{1}$ etc.
@@
 C.2 Correctness and running time
-Both are straightforward adaptation of what was done for Annotated Dominated Cluster Deletion. For example Section B.3 also exactly explain the running time of the current algorithm.
+Both are straightforward adaptation of what was done for Annotated Dominated Cluster Deletion. For example, Section˜B.3 also exactly explains the running time of the current algorithm.
 For correctness, pick a realized $(k,d)$-extended-mark $M_{x}$ of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$.